### v5

# Non-Equilibrium Steady States and Quantum Chaos in a three-site Driven-Dissipative Bose-Hubbard Chains base on Self-Consistent Mean-Field Approach

Abstract
We investigate the non-equilibrium dynamics and steady-state properties of a driven-dissipative Bose-Hubbard chain using a self-consistent Gutzwiller mean-field (GMF) approach. By employing a robust Picard iteration scheme, we solve the non-linear master equation for the non-equilibrium steady state (NESS) in the presence of strong Kerr nonlinearity. We identify two distinct dynamical regimes governed by the interplay between coherent drive, dissipation, and interaction: a regular quasilinear regime and a chaotic regime. Linear stability analysis reveals that the transition to the chaotic regime is triggered by parametric instabilities arising from the drive-induced coherence. Furthermore, we characterize the onset of quantum chaos by calculating the out-of-time-order correlator (OTOC). Our results show that in the strong coupling regime, the OTOC exhibits rapid exponential growth and saturation, providing a clear signature of information scrambling in this open quantum system. The proposed numerical framework offers an efficient pathway to explore many-body correlations in larger photonic lattices.
I Introduction
Driven-dissipative quantum many-body systems have emerged as a versatile platform for exploring novel phases of matter that have no equilibrium counterparts. Among these, the driven-dissipative Bose-Hubbard model, realizable in superconducting circuit QED and coupled cavity arrays, serves as a paradigm for studying the competition between coherent drive, particle loss, and on-site nonlinearity. Unlike closed systems governed by thermodynamics, these open systems settle into a NESS determined by the balance of gain and loss. A central open question in this field is characterizing the transition from regular, mean-field-like dynamics to complex, chaotic behavior where quantum correlations play a dominant role.In this work, we address this problem by analyzing a one-dimensional Bose-Hubbard chain under coherent driving. We develop a numerical approach based on the Gutzwiller Mean-Field approximation, utilizing a Picard iteration technique to rigorously solve the non-linear self-consistency equations for the NESS. This method allows us to treat the on-site Kerr nonlinearity exactly, capturing high-order local correlations $\langle\hat{n}\hat{a}\rangle$ that are crucial for determining stability. We derive the effective Liouvillian governing the fluctuations around the NESS and demonstrate that the system undergoes a dynamical phase transition characterized by the closure of the Liouvillian gap. To explicitly diagnose the quantum chaotic nature of the high-drive phase, we compute the OTOC. We observe that in the strong coupling regime, the OTOC saturates to a macroscopic value following an initial period of exponential growth, signaling the scrambling of quantum information. Our findings establish a clear link between parametric instability in the mean-field spectrum and the emergence of many-body chaos in open quantum systems.
II model
We consider the system involving the interaction between local boson and nonlocal boson, which reads
$H_{b}=\sum_{r=1}^{L}\left(\Delta\hat{a}^{\dagger}_{r}\hat{a}_{r}+\frac{1}{2}U\hat{a}^{\dagger}_{r}\hat{a}^{\dagger}_{r}\hat{a}_{r}\hat{a}_{r}\right)-\sum_{r=1}^{L-1}J\left(\hat{a}^{\dagger}_{l+1}\hat{a}_{r}+\hat{a}^{\dagger}_{r}\hat{a}_{l+1}\right)+F(\hat{a}^{\dagger}_{1}+\hat{a}_{1})$ (1)
where $\Delta$ is the pump-to-resonator detuning between a drive frequency and the resonator’s natural frequency ($\Delta=\omega_{d}-\omega_{0}$). $\sum_{r=1}^{L}\frac{1}{2}U\hat{a}^{\dagger}_{r}\hat{a}^{\dagger}_{r}\hat{a}_{r}\hat{a}_{r}=\frac{U}{2}\sum_{r=1}^{L}\hat{n}_{r}(\hat{n}_{r}-1)$ is the term describe the on-site interaction. $U$ is the strength of the onsite Kerr nonlinearity. $J$ is the hopping amplitude. $F$ is the strength (amplitude) of coherent drive, which describes an external driving field that coherently pumps energy into the chain by creating ($\hat{a}^{\dagger}_{1}$) and destroying ($\hat{a}_{1}$) bosons. For $F>0$ and $\gamma>0$, the drive continuously pumps energy in and the loss continuously drains them. The system settles into a stable state with a net flow of particles, which is the definition of a non-equilibrium (and non-thermal) state. As a result, it reach a unique NESS at long-time limit.
III Gutzwiller mean-field approach
Using the Gutzwiller mean-field decoupling, we decompose the bath boson operator into the mean field part and fluctuational part, $a_{l}=\psi_{l}(t)+\delta a_{l}(t)$, where $\psi_{r}(t)=\langle\hat{b}_{r}\rangle$ is the classical complex condensate amplitude, and $\delta a_{l}(t)$ is the quantum fluctuation operator. Then the hopping interaction term becomes $\hat{a}_{r}^{\dagger}\hat{a}_{r^{\prime}}=(\psi_{r}^{*}+\delta\hat{a}_{r}^{\dagger})(\psi_{r^{\prime}}+\delta\hat{a}_{r^{\prime}})=\psi_{r}^{*}\psi_{r^{\prime}}+\psi_{r}^{*}\delta\hat{a}_{r^{\prime}}+\delta\hat{a}_{r}^{\dagger}\psi_{r^{\prime}}+\delta\hat{a}_{r}^{\dagger}\delta\hat{a}_{r^{\prime}}$ where the quantum fluctuations described in last term is neglected in mean-field approximation.
For any site $r$ that is not the impurity site, the dynamics are governed by a single-site master equation
$\frac{d\hat{\rho}_{r}}{dt}=-i[H_{\text{MF},r},\hat{\rho}_{r}]+\mathcal{D}_{r}[\hat{\rho}_{r}]$ (2)
The mean-field Hamiltonian for this site is
$H_{\text{MF},r}=\Delta\hat{a}_{r}^{\dagger}\hat{a}_{r}+\frac{U}{2}\hat{a}_{r}^{\dagger}\hat{a}_{r}^{\dagger}\hat{a}_{r}\hat{a}_{r}-J\left(\hat{a}^{\dagger}_{r}\hat{a}_{r-1}+\hat{a}^{\dagger}_{r-1}\hat{a}_{r}\right)-J\left(\hat{a}^{\dagger}_{r+1}\hat{a}_{r}+\hat{a}^{\dagger}_{r}\hat{a}_{r+1}\right)$ (3)
$=\Delta\hat{a}_{r}^{\dagger}\hat{a}_{r}+\frac{U}{2}\hat{a}_{r}^{\dagger}\hat{a}_{r}^{\dagger}\hat{a}_{r}\hat{a}_{r}-J\left[\left(\Psi_{r-1}+\Psi_{r+1}\right)\hat{a}_{r}^{\dagger}+\left(\Psi_{r-1}^{*}+\Psi_{r+1}^{*}\right)\hat{a}_{r}\right]$
where we treat the neighbors (r1r-1 and $r+1$) as environment. This is the single-site Hamiltonian where the quantum operators $\hat{a}_{l}$ and $\hat{a}_{l}^{\dagger}$ interact with a classical, external field $\left(\Psi_{l-1}+\Psi_{l+1}\right)$ that represents the average influence of all its neighbors. where the operator $\hat{a}_{l}^{\dagger}$ is coupled to the classical field $\Psi_{l-1}+\Psi_{l+1}$.
In Gutzwiller approximation, the effective single-site Hamiltonian for a "bath" site $r(\neq j)$
$H_{\text{MF},r}=$ $\Delta\hat{a}_{r}^{\dagger}\hat{a}_{r}+\frac{U}{2}\hat{a}_{r}^{\dagger}\hat{a}_{r}^{\dagger}\hat{a}_{r}\hat{a}_{r}$ (4)
$-J\left[\left((1-\delta_{r,1})\Psi_{r-1}+(1-\delta_{r,L})\Psi_{r+1}\right)\hat{a}_{r}^{\dagger}+\left((1-\delta_{r,1})\Psi_{r-1}^{*}+(1-\delta_{r,L})\Psi_{r+1}^{*}\right)\hat{a}_{r}\right]+F(a_{1}^{{\dagger}}+a_{1})\delta_{r,1}.$
The Gutzwiller mean-field approximation simplifies this by factorizing the total state. It assumes there are no quantum correlations (entanglement) between different sites. The total density matrix is approximated as a simple product of independent, single-site density matrices, $\hat{\rho}_{\text{total}}(t)\approx\hat{\rho}_{1}(t)\otimes\hat{\rho}_{2}(t)\otimes\dots\otimes\hat{\rho}_{j}(t)\otimes\dots\otimes\hat{\rho}_{L}(t)$. Note that the vacuum state $|0\rangle_{r}$ is an eigenstate of the local Hamiltonian $H_{loc,r}:=\Delta\hat{a}_{r}^{\dagger}\hat{a}_{r}+\frac{U}{2}\hat{a}_{r}^{\dagger}\hat{a}_{r}^{\dagger}\hat{a}_{r}\hat{a}_{r}$ with also an eigenvalue of 0, $|0\rangle_{r}$ is also an eigenstate of $a_{r}^{\dagger}a_{r}$ with an eigenvalue of 0. Thus we have $H_{loc,r}|0\rangle_{r}=a_{r}|0\rangle_{r}=a_{r}^{\dagger}a_{r}|0\rangle_{r}=0$, $a_{r}^{\dagger}a_{r}|n\rangle_{r}=n|n\rangle_{r}$ ($n\geq 0$), $a_{r}|0\rangle_{r}^{\prime}=|0\rangle_{r}^{\prime}$ for $r\neq r^{\prime}$.
To obtain the self-consistent solution of NESS, we assume the chain starts in vacuum state $\rho_{r}(0)=|0\rangle_{r}\langle 0|_{r}$, whose expectation is $\Psi_{r}(0)={\rm Tr}[a_{r}\rho_{r}(0)]=0$. For the drive site $r=1$, $\rho_{1}(0)=|0\rangle_{1}\langle 0|_{1}$,
$\mathcal{L}_{1}(\rho_{1}(0))=-i[H_{{\rm MF},1}(0),\rho_{1}(0)]+\mathcal{D}_{1}[\rho_{1}(0)]$ (5)
$=-i[\Delta\hat{a}_{1}^{\dagger}\hat{a}_{1}+\frac{U}{2}a_{1}^{\dagger}\hat{a}_{1}^{\dagger}a_{1}a_{1}-J(\Psi_{2}(0)a_{r}^{\dagger}+\Psi_{2}^{*}(0)a_{r})+F(a_{1}^{{\dagger}}+a_{1}),\rho_{1}(0)]$
$=-iF\left((\hat{a}^{\dagger}_{1}+\hat{a}_{1})|0\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 0|_{1}(a^{\dagger}_{1}+a_{1})\right)$
$=-iF\left(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1}\right)$
where we use $\mathcal{D}_{1}[\rho_{1}(0)]=0$, $\Delta\hat{a}_{1}^{\dagger}\hat{a}_{1}+\frac{U}{2}a_{1}^{\dagger}\hat{a}_{1}^{\dagger}a_{1}a_{1},\rho_{1}(0)]=0$, $a_{1}|0\rangle_{1}=0$, $a_{1}|1\rangle_{1}=|0\rangle_{1}$, $a^{\dagger}_{1}|0\rangle_{1}=|1\rangle_{1}$, $a^{\dagger}_{1}|1\rangle_{1}=\sqrt{2}|2\rangle_{1}$, $\langle 0|_{1}a^{{\dagger}}_{1}=(a_{1}|0\rangle_{1})^{{\dagger}}=0$, $\langle 1|_{1}a^{{\dagger}}_{1}=(a_{1}|1\rangle_{1})^{{\dagger}}=\langle 0|$, $\langle 1|_{1}a_{1}=(a^{{\dagger}}_{1}|1\rangle_{1})^{{\dagger}}=\sqrt{2}\langle 2|$. For number (Fock) basis $\{|k\rangle\}$, ${\rm Tr}[a|n\rangle\langle m|]=\sum_{k=0}^{\infty}\langle k|a|n\rangle\langle m|k\rangle=\sum_{k=0}^{\infty}\sqrt{n}\delta_{k,n-1}\delta_{m,k}=\sqrt{n}\delta_{m,n-1}$. Similarly, $\text{Tr}[\hat{a}_{l}\sum_{n,m}c_{nm}|n\rangle\langle m|]=\sum_{n,m}c_{nm}\text{Tr}[\hat{a}_{l}|n\rangle\langle m|]=\sum_{n,m}c_{nm}(\sqrt{n}\delta_{m,n-1})=\sum_{n=1}^{\infty}c_{n,n-1}\sqrt{n}$. Thus
$\rho_{1}(dt)=\rho_{1}(0)+\mathcal{L}_{1}(\rho_{1}(0))dt=|0\rangle_{1}\langle 0|_{1}-iFdt\left(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1}\right)$ (6)
where we use the Euler method approximation
$\rho_{r}(dt)=\rho_{r}(0)+\left(\frac{d\rho_{r}}{dt}\bigg|_{t=0}\right)dt=\rho_{r}(0)+\mathcal{L}_{l}(\rho_{r}(0))dt.$ (7)
The corresponding condensate amplitude reads
$\Psi_{1}(dt)=\text{Tr}\left[\hat{a}_{1}\cdot\left(|0\rangle_{1}\langle 0|_{1}-iFdt(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})\right)\right]=\text{Tr}[-iFdt|0\rangle_{1}\langle 0|_{1}]=-iFdt$ (8)
For bulk site $1<r<L$, $\rho_{r}(0)=|0\rangle_{r}\langle 0|_{r}$, and $\mathcal{L}_{r}(\rho_{r}(0))=-i[\Delta\hat{a}_{r}^{\dagger}\hat{a}_{r}+\frac{U}{2}a_{r}^{\dagger}\hat{a}_{r}^{\dagger}a_{r}a_{r},\rho_{r}(0)]+\mathcal{D}_{r}[\rho_{r}(0)]=0$ since there is no loss in bulk $\mathcal{D}_{r}[\rho_{r}(0)]=0$. Thus $\rho_{r}(dt)=|0\rangle_{r}\langle 0|_{r}$ and $\Psi_{r}(dt)={\rm Tr}[a_{r}\rho_{r}(dt)]=0$, i.e., it remains in vacuum state in the first time step. For the loss site we have
$\mathcal{L}_{L}(\rho_{L}(0))=-i[H_{{\rm MF},L}(0),\rho_{L}(0)]+\mathcal{D}_{L}[\rho_{L}(0)]=\mathcal{D}_{L}[\rho_{L}(0)]$ (9)
$=\gamma\hat{a}_{L}|0\rangle_{L}\langle 0|_{L}\hat{a}_{L}^{\dagger}-\frac{\gamma}{2}\{\hat{a}_{L}^{\dagger}\hat{a}_{L},|0\rangle_{L}\langle 0|_{L}\}=0$
where we use $a_{L}|0\rangle=0$ and $a_{L}^{\dagger}a_{L}|0\rangle=0$. Thus $\rho_{L}(dt)=|0\rangle_{L}\langle 0|_{L}$ and $\Psi_{L}(dt)=0$. Now the mean-field single-site Hamiltonian for bath reads
$H_{{\rm MF},1}(dt)=H_{loc,1}+F(\hat{a}^{\dagger}_{1}+\hat{a}_{1}),$ (10)
$H_{{\rm MF},2}(dt)=H_{loc,2}-J[(\Psi_{1}(dt)+\Psi_{3}(dt))a_{2}^{\dagger}+h.c.]=H_{loc,2}-J[(-iFdt)a_{2}^{\dagger}+(iFdt)a_{2}],$
$H_{{\rm MF},r}(dt)=H_{loc,r},(for3\leq r\leq L).$
At $2dt$, the densities read
$\rho_{1}(2dt)=\rho_{1}(dt)+\mathcal{L}_{1}(\rho_{1}(dt))dt$ (11)
$=|0\rangle_{1}\langle 0|_{1}-iFdt\left(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1}\right)-i[H_{{\rm MF},1}(dt),\rho_{1}(dt)]dt+\mathcal{D}_{1}[\rho_{1}(dt)]dt$
$=|0\rangle_{1}\langle 0|_{1}-iFdt\left(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1}\right)-i[H_{loc,1}+F(a_{1}^{{\dagger}}+a_{1}),\rho_{1}(dt)]dt+\mathcal{D}_{1}[\rho_{1}(dt)]dt$
$=|0\rangle_{1}\langle 0|_{1}-2iFdt(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})-F\Delta(dt)^{2}(|1\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 1|_{1})$
$-F^{2}(dt)^{2}\left[2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1})\right],$
$\rho_{2}(2dt)=\rho_{2}(dt)+\mathcal{L}_{2}(\rho_{2}(dt))dt=|0\rangle_{2}\langle 0|_{2}-i[H_{{\rm MF},2}(dt),\rho_{2}(dt)]dt+\mathcal{D}_{2}[\rho_{2}(dt)]dt$
$=|0\rangle_{2}\langle 0|_{2}-i[H_{loc,2}-J[(-iFdt)a_{2}^{\dagger}+(iFdt)a_{2}],\rho_{2}(dt)]dt$
$=|0\rangle_{2}\langle 0|_{2}+iJ\left[((-iFdt)\hat{a}_{2}^{\dagger}+(iFdt)\hat{a}_{2}),|0\rangle\langle 0|\right]dt$
$=|0\rangle_{2}\langle 0|_{2}+iJ(-iFdt)(|1\rangle_{2}\langle 0|_{2}+|0\rangle_{2}\langle 1|_{2})dt$
$=|0\rangle_{2}\langle 0|_{2}+JF(dt)^{2}(|1\rangle_{2}\langle 0|_{2}+|0\rangle_{2}\langle 1|_{2}),$
$\rho_{3}(2dt)=\rho_{3}(dt)+\mathcal{L}_{3}(\rho_{3}(dt))dt=|0\rangle_{3}\langle 0|_{3}-i[H_{loc,3}-J[(\Psi_{2}(dt)+\Psi_{4}(dt))\hat{a}_{3}^{\dagger}+h.c.],\rho_{3}(dt)]dt+\mathcal{D}_{3}[\rho_{3}(dt)]dt$
$=|0\rangle_{3}\langle 0|_{3}-i[H_{loc,3},\rho_{3}(dt)]dt+\mathcal{D}_{3}[\rho_{3}(dt)]dt=|0\rangle_{3}\langle 0|_{3},$
$\cdots,$
$\rho_{L}(2dt)\approx|0\rangle_{L}\langle 0|_{L},$
where the dissipation rate $\gamma_{1}=\gamma_{L}=0$. We use the following relations
$H_{loc,r}|0\rangle_{r}=\langle 0|_{r}H_{loc,r}=0,\ H_{loc,r}|1\rangle_{r}=\Delta|1\rangle_{r},\ \langle 1|_{r}H_{loc,r}=\Delta\langle 1|_{r},H_{loc,r}|2\rangle_{r}=(2\Delta+U)|2\rangle_{r},$ (12)
$[H_{loc,1},|0\rangle_{1}\langle 0|_{1}]=0,[F(a_{1}^{{\dagger}}+a_{1}),|0\rangle_{1}\langle 0|_{1}]=F(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1}),$
$[H_{loc,1},-iFdt(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})]=-iFdt\Delta(|1\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 1|_{1}),$
$[F(\hat{a}^{\dagger}_{1}+\hat{a}_{1}),-iFdt(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})]=-iF^{2}dt\left[2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}\left(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1}\right)\right],$
$[H_{loc,1}+F(a_{1}^{{\dagger}}+a_{1}),\rho_{1}(dt)]=F(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})-iFdt\Delta(|1\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 1|_{1})$
$-iF^{2}dt\left[2|0\rangle\langle 0|-2|1\rangle\langle 1|+\sqrt{2}\left(|2\rangle\langle 0|+|0\rangle\langle 2|\right)\right],$
$H_{loc,r}|n\rangle_{r}=\Delta n|n\rangle_{r}+\frac{U}{2}n(n-1)|n\rangle_{r},$
$\langle n|_{r}H_{loc,r}=\Delta n\langle n|_{r}+\frac{U}{2}n(n-1)\langle n|_{r},$
Note that even for $\gamma_{1}\neq 0$,
$\mathcal{D}_{1}[\rho_{1}(dt)]=\gamma_{1}a_{1}\rho_{1}(dt)a^{{\dagger}}_{1}-\frac{\gamma_{1}}{2}\{a_{1}^{{\dagger}}a_{1},\rho_{1}(dt)\}$ (13)
$=-iFdt\gamma_{1}|0\rangle\langle 1|-\frac{\gamma_{1}}{2}\{a_{1}^{{\dagger}}a_{1},-iFdt\left(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1}\right)\}=-iFdt\gamma_{1}|0\rangle\langle 1|+\frac{iFdt\gamma_{1}}{2}\left(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1}\right),$
$\mathcal{D}_{2}[\rho_{2}(dt)]=\gamma_{2}a_{2}|0\rangle_{2}\langle 0|_{2}a^{{\dagger}}_{2}-\frac{\gamma_{2}}{2}\{a_{2}^{{\dagger}}a_{2},|0\rangle_{2}\langle 0|_{2}\}=0.$
where $\{a_{1}^{\dagger}a_{1},|1\rangle\langle 0|\}=|1\rangle\langle 0|$, $\{a^{{\dagger}}_{1}a_{1},|0\rangle_{1}\langle 0|_{1}\}=|0\rangle_{1}\langle 0|_{1}$.
At $2dt$, the mean-field Hamiltonian read
$H_{MF,1}(2dt)=H_{loc,1}+F(a^{\dagger}_{1}+a_{1})-J[\Psi_{2}(2dt)a_{1}^{\dagger}+h.c.]$ (14)
$=H_{loc,1}+F(a^{\dagger}_{1}+a_{1})-J[JF(dt)^{2}a_{1}^{\dagger}+h.c.]$
$=H_{loc,1}+F(a^{\dagger}_{1}+a_{1})-J^{2}F(dt)^{2}(a_{1}^{\dagger}+a_{1}),$
$H_{MF,2}(2dt)=H_{loc,2}-J[(\Psi_{1}(2dt)+\Psi_{3}(2dt))a_{2}^{\dagger}+h.c.]$
$=H_{loc,2}-J[(-2iFdt-F\Delta(dt)^{2})a_{2}^{\dagger}+h.c.],$
$H_{MF,3}(2dt)=H_{loc,3}-J\left[(\Psi_{2}(2dt)+\Psi_{4}(2dt))a_{3}^{\dagger}+h.c.\right]=H_{loc,3}-J^{2}F(dt)^{2}(a_{3}^{\dagger}+a_{3})$
where the amplitudes read $\Psi_{1}(2dt)=-2iFdt-F\Delta(dt)^{2}$, $\Psi_{2}(2dt)=JF(dt)^{2}$, $\Psi_{3}(2dt)=0$, $\cdots$.
At time $2dt$, we have
$\mathcal{D}_{1}[\hat{\rho}_{1}(2dt)]=$ $\mathcal{D}_{1}[|0\rangle_{1}\langle 0|_{1}-2iFdt(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})-F\Delta(dt)^{2}(|1\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 1|_{1})$ (15)
$-F^{2}(dt)^{2}\left[2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1})\right]]$
$=$ $-2iFdt(-\frac{\gamma_{1}}{2}|1\rangle_{1}\langle 0|_{1}+\frac{\gamma_{1}}{2}|0\rangle_{1}\langle 1|_{1})$
$+\frac{F\Delta\gamma_{1}}{2}(dt)^{2}(|1\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 1|_{1})$
$+F^{2}\gamma_{1}(dt)^{2}\left[2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1})\right],$
$[H_{MF,1}(2dt),\rho_{1}(2dt)]=$ $[H_{loc,1}+F(a^{\dagger}_{1}+a_{1})-J^{2}F(dt)^{2}(a_{1}^{\dagger}+a_{1}),|0\rangle_{1}\langle 0|_{1}-2iFdt(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})$
$-F\Delta(dt)^{2}(|1\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 1|_{1})$
$-F^{2}(dt)^{2}\left[2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1})\right]]$
$=(F-J^{2}F(dt)^{2})(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})-2iF\Delta dt(|1\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 1|_{1})$
$-2iF^{2}dt(2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1}))$
$-F\Delta^{2}(dt)^{2}(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})-4F^{3}(dt)^{2}(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})$
$-\sqrt{2}F^{2}(3\Delta+U)(dt)^{2}(|2\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 2|_{1})$
$-6F^{3}(dt)^{2}(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})+O(F^{3}(dt)^{2})$
$+2iF^{2}J^{2}(dt)^{3}(2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1})),$
$\mathcal{D}_{2}[\rho_{2}(2dt)]$ $=JF(dt)^{2}\left(-\frac{\gamma_{2}}{2}|1\rangle_{2}\langle 0|_{2}-\frac{\gamma_{2}}{2}|0\rangle_{2}\langle 1|_{2}\right),$
$[H_{MF,2}(2dt),\rho_{2}(2dt)]=$ $\qquad 2iJFdt(|1\rangle_{2}\langle 0|_{2}+|0\rangle_{2}\langle 1|_{2})$
$+2JF\Delta(dt)^{2}(|1\rangle_{2}\langle 0|_{2}-|0\rangle_{2}\langle 1|_{2})$
$+2iJ^{2}F^{2}(dt)^{3}(\sqrt{2}|2\rangle_{2}\langle 0|_{2}+|1\rangle_{2}\langle 1|_{2}-\sqrt{2}|0\rangle_{2}\langle 2|_{2})+O(dt^{4}),$
$\mathcal{D}_{3}[\rho_{3}(2dt)]$ $=\mathcal{D}_{3}[|0\rangle_{3}\langle 0|_{3}]=0,$
$[H_{MF,3}(2dt),\rho_{3}(2dt)]$ $=-J^{2}F(dt)^{2}(|1\rangle_{3}\langle 0|_{3}-|0\rangle_{3}\langle 1|_{3}),$
where we use the following results
$[H_{loc,2},|0\rangle_{2}\langle 0|_{2}]=0,$ (16)
$[-J(-2iFdt-F\Delta(dt)^{2})\hat{a}_{2}^{\dagger}+h.c.,|0\rangle_{2}\langle 0|_{2}]=J(2iFdt+F\Delta(dt)^{2})|1\rangle_{2}\langle 0|_{2}+J(2iFdt-F\Delta(dt)^{2})|0\rangle_{2}\langle 1|_{2},$
$[H_{loc,2},JF(dt)^{2}(|1\rangle_{2}\langle 0|_{2}+|0\rangle_{2}\langle 1|_{2})]=JF(dt)^{2}\Delta(|1\rangle_{2}\langle 0|_{2}-|0\rangle_{2}\langle 1|_{2}),$
$[-J(-2iFdt-F\Delta(dt)^{2})\hat{a}_{2}^{\dagger}+h.c.,JF(dt)^{2}(|1\rangle_{2}\langle 0|_{2}+|0\rangle_{2}\langle 1|_{2})]=$
$J(2iFdt+F\Delta(dt)^{2})|1\rangle_{2}\langle 0|_{2}+J(2iFdt-F\Delta(dt)^{2})|0\rangle_{2}\langle 1|_{2}$
$+JF\Delta(dt)^{2}(|1\rangle_{2}\langle 0|_{2}-|0\rangle_{2}\langle 1|_{2})+2iJ^{2}F^{2}(dt)^{3}(\sqrt{2}|2\rangle_{2}\langle 0|_{2}+|1\rangle_{2}\langle 1|_{2}-\sqrt{2}|0\rangle_{2}\langle 2|_{2})+O(dt^{4}).$
Then we can obtain the density and the related amplitude as
$\rho_{1}(3dt)$ $=\rho_{1}(2dt)+\mathcal{L}_{1}(\rho_{1}(2dt))dt=\rho_{1}(2dt)+\left(-i[H_{MF,1}(2dt),\rho_{1}(2dt)]+\mathcal{D}_{1}[\rho_{1}(2dt)]\right)dt$ (17)
$=|0\rangle_{1}\langle 0|_{1}-3iFdt(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})$
$+\left(-3F\Delta(dt)^{2}+iF\gamma_{1}(dt)^{2}\right)|1\rangle_{1}\langle 0|_{1}+\left(-3F\Delta(dt)^{2}-iF\gamma_{1}(dt)^{2}\right)|0\rangle_{1}\langle 1|_{1}$
$-3F^{2}(dt)^{2}\left[2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1})\right]$
$+\left(iF\Delta^{2}+6iF^{3}+iJ^{2}F+\frac{\gamma_{1}F\Delta}{2}\right)(dt)^{3}|1\rangle_{1}\langle 0|_{1}$
$+\left(-iF\Delta^{2}-6iF^{3}-iJ^{2}F+\frac{\gamma_{1}F\Delta}{2}\right)(dt)^{3}|0\rangle_{1}\langle 1|_{1}$
$+i\sqrt{2}F^{2}(3\Delta+U)(dt)^{3}(|2\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 2|_{1})$
$+\gamma_{1}F^{2}(dt)^{3}(2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1})),$
$\Psi_{1}(3dt)$ $=\text{Tr}[a_{1}\rho_{1}(3dt)]=-3iFdt+(-3F\Delta+iF\gamma_{1})(dt)^{2}$
$+\left(i(J^{2}F+F\Delta^{2}+6F^{3})+\frac{\gamma_{1}F\Delta}{2}\right)(dt)^{3},$
$\rho_{2}(3dt)$ $=\rho_{2}(2dt)+\mathcal{L}_{2}(\rho_{2}(2dt))dt=\rho_{2}(2dt)+\left(-i[H_{MF,2}(2dt),\rho_{2}(2dt)]+\mathcal{D}_{2}[\rho_{2}(2dt)]\right)dt$
$=|0\rangle_{2}\langle 0|_{2}+3JF(dt)^{2}(|1\rangle_{2}\langle 0|_{2}+|0\rangle_{2}\langle 1|_{2})$
$+\left(-2iJF\Delta-\frac{\gamma_{2}JF}{2}\right)(dt)^{3}|1\rangle_{2}\langle 0|_{2}+\left(2iJF\Delta-\frac{\gamma_{2}JF}{2}\right)(dt)^{3}|0\rangle_{2}\langle 1|_{2},$
$\Psi_{2}(3dt)$ $=\text{Tr}[a_{2}\rho_{2}(3dt)]=3JF(dt)^{2}+\left(-2iJF\Delta-\frac{\gamma_{2}JF}{2}\right)(dt)^{3},$
$\rho_{3}(3dt)$ $=\rho_{3}(2dt)+\mathcal{L}_{3}(\rho_{3}(2dt))dt=|0\rangle_{3}\langle 0|_{3}-i[H_{loc,3}-J^{2}F(dt)^{2}(a_{3}^{\dagger}+a_{3}),|0\rangle_{3}\langle 0|_{3}]dt$
$=|0\rangle_{3}\langle 0|_{3}+iJ^{2}F(dt)^{2}(|1\rangle_{3}\langle 0|_{3}-|0\rangle_{3}\langle 1|_{3})dt,$
$\Psi_{3}(3dt)$ $=\text{Tr}[a_{3}\rho_{3}(3dt)]=iJ^{2}F(dt)^{3}.$
Thus $\Psi_{3}$ is non-zero since $3dt$, and will next activate the Hamiltonian $H_{MF,4}$ and create a non-zero $\Psi_{4}$.
At time $3dt$, the mean-field Hamiltonians are
$H_{MF,1}(3dt)=H_{loc,1}+F(\hat{a}^{\dagger}_{1}+\hat{a}_{1})-J[\Psi_{2}(3dt)\hat{a}_{1}^{\dagger}+h.c.],$ (18)
$H_{MF,2}(3dt)=H_{loc,2}-J[(\Psi_{1}(3dt)+\Psi_{3}(3dt))\hat{a}_{2}^{\dagger}+h.c.],$
$H_{MF,3}(3dt)=H_{loc,3}-J[\Psi_{3}(3dt)\hat{a}_{3}^{\dagger}+h.c.]$
the corresponding dissipators and commutators read
$\mathcal{D}_{1}[\rho_{1}(3dt)]=i\frac{3\gamma_{1}F}{2}dt(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})+\left(\frac{3\gamma_{1}F\Delta}{2}-\frac{i\gamma_{1}^{2}F}{2}\right)(dt)^{2}|1\rangle_{1}\langle 0|_{1}$ (19)
$+\left(\frac{3\gamma_{1}F\Delta}{2}+\frac{i\gamma_{1}^{2}F}{2}\right)(dt)^{2}|0\rangle_{1}\langle 1|_{1}$
$+3\gamma_{1}F^{2}(dt)^{2}\left[2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1})\right]$
$+\left(-\frac{i\gamma_{1}}{2}(F\Delta^{2}+6F^{3}+J^{2}F)-\frac{\gamma_{1}^{2}F\Delta}{4}\right)(dt)^{3}|1\rangle_{1}\langle 0|_{1}+\left(\frac{i\gamma_{1}}{2}(F\Delta^{2}+6F^{3}+J^{2}F)-\frac{\gamma_{1}^{2}F\Delta}{4}\right)(dt)^{3}|0\rangle_{1}\langle 1|_{1}$
$+\left(-i\sqrt{2}\gamma_{1}F^{2}(3\Delta+U)-\sqrt{2}\gamma_{1}^{2}F^{2}\right)(dt)^{3}|2\rangle_{1}\langle 0|_{1}$
$+\left(i\sqrt{2}\gamma_{1}F^{2}(3\Delta+U)-\sqrt{2}\gamma_{1}^{2}F^{2}\right)(dt)^{3}|0\rangle_{1}\langle 2|_{1}-2\gamma_{1}^{2}F^{2}(dt)^{3}(|0\rangle_{1}\langle 0|_{1}-|1\rangle_{1}\langle 1|_{1}),$
$[H_{MF,1}(3dt),\rho_{1}(3dt)]=F(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})-3iF\Delta dt(|1\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 1|_{1})$
$-3iF^{2}dt\left(2|0\rangle_{1}\langle 0|_{1}-2|1\rangle_{1}\langle 1|_{1}+\sqrt{2}(|2\rangle_{1}\langle 0|_{1}+|0\rangle_{1}\langle 2|_{1})\right)$
$-3J^{2}F(dt)^{2}(|1\rangle_{1}\langle 0|_{1}-|0\rangle_{1}\langle 1|_{1})-J\left(-2iJF\Delta-\frac{\gamma_{2}JF}{2}\right)(dt)^{3}|1\rangle_{1}\langle 0|_{1}+J\left(2iJF\Delta-\frac{\gamma_{2}JF}{2}\right)(dt)^{3}|0\rangle_{1}\langle 1|_{1},$
$\mathcal{D}_{2}[\rho_{2}(3dt)]=-\frac{3\gamma_{2}JF}{2}(dt)^{2}(|1\rangle_{2}\langle 0|_{2}+|0\rangle_{2}\langle 1|_{2})$
$+\left(i\gamma_{2}JF\Delta+\frac{\gamma_{2}^{2}JF}{4}\right)(dt)^{3}|1\rangle_{2}\langle 0|_{2}+\left(-i\gamma_{2}JF\Delta+\frac{\gamma_{2}^{2}JF}{4}\right)(dt)^{3}|0\rangle_{2}\langle 1|_{2},$
$[H_{MF,2}(3dt),\rho_{2}(3dt)]=3iJFdt(|1\rangle_{2}\langle 0|_{2}+|0\rangle_{2}\langle 1|_{2})+(6JF\Delta-iJF\gamma_{1})(dt)^{2}|1\rangle_{2}\langle 0|_{2}$
$+(-6JF\Delta-iJF\gamma_{1})(dt)^{2}|0\rangle_{2}\langle 1|_{2}+J(i(2J^{2}F+F\Delta^{2}+6F^{3})+\frac{\gamma_{1}F\Delta}{2})(dt)^{3}|1\rangle\langle 0|$
$+J(-i(2J^{2}F+F\Delta^{2}+6F^{3})+\frac{\gamma_{1}F\Delta}{2})(dt)^{3}|0\rangle\langle 1|,$
$\mathcal{D}_{3}[\rho_{3}(3dt)]=-i\frac{\gamma_{3}J^{2}F}{2}(dt)^{3}(|1\rangle_{3}\langle 0|_{3}-|0\rangle_{3}\langle 1|_{3}),$
$[H_{MF,3}(3dt),\rho_{3}(3dt)]=[-J(3JF(dt)^{2}\hat{a}_{3}^{\dagger}+\text{H.c.}),|0\rangle\langle 0|]$
$=-3J^{2}F(dt)^{2}(|1\rangle_{3}\langle 0|_{3}-|0\rangle_{3}\langle 1|_{3})+J(-2iJF\Delta-\frac{\gamma_{2}JF}{2})(dt)^{3}|1\rangle_{3}\langle 0|_{3}-J(2iJF\Delta-\frac{\gamma_{2}JF}{2})(dt)^{3}|0\rangle_{3}\langle 1|_{3},$
and we can then obtain the amplitudes of the next time step
$\Psi_{1}(4dt)=-4iFdt+(-6F\Delta+i\frac{5\gamma_{1}F}{2})(dt)^{2}+\left(i(4J^{2}F+4F\Delta^{2}+12F^{3})+3\gamma_{1}F\Delta-\frac{i\gamma_{1}^{2}F}{2}\right)(dt)^{3},$ (20)
$\Psi_{2}(4dt)=6JF(dt)^{2}+\left(-8iJF\Delta-JF\gamma_{1}-2\gamma_{2}JF\right)(dt)^{3},$
$\Psi_{3}(4dt)=4iJ^{2}F(dt)^{3}.$
We conclude in Table.I the amplitudes for until the time step $4dt$ for the three sites. The leading term of amplitudes exhibits $\Psi_{N}(Ndt)\sim O((dt)^{N})$. The coefficients of these leading terms grow polynomially with each time step. Keeping only the leading term, we have $\Psi_{1}(Ndt)=-iNFdt$, $\Psi_{2}(Ndt)=\frac{N(N-1)}{2}JF(dt)^{2}$, $\Psi_{3}(Ndt)=\frac{N(N-1)(N-2)}{6}iJ^{2}F(dt)^{3}$.
Table 1: Perturbative Expansion of Mean-Field Amplitudes $\Psi_{l}(N\cdot dt)$
Time Step (N) $\Psi_{1}(N\cdot dt)$ $\Psi_{2}(N\cdot dt)$ $\Psi_{3}(N\cdot dt)$
1 iFdt-iFdt 0 0
2 $2iFdt-F\Delta(dt)^{2}$ $JF(dt)^{2}$ 0
3
$3iFdt+(-3F\Delta+iF\gamma_{1})(dt)^{2}$
$\left(i(J^{2}F+F\Delta^{2}+6F^{3})+\frac{\gamma_{1}F\Delta}{2}\right)(dt)^{3}$
$3JF(dt)^{2}$
$\left(-2iJF\Delta-\frac{\gamma_{2}JF}{2}\right)(dt)^{3}$
$iJ^{2}F(dt)^{3}$
4
$4iFdt+(-6F\Delta+i\frac{5\gamma_{1}F}{2})(dt)^{2}$
$\left(i(4J^{2}F+4F\Delta^{2}+12F^{3})+3\gamma_{1}F\Delta-\frac{i\gamma_{1}^{2}F}{2}\right)(dt)^{3}$
$6JF(dt)^{2}$
$\left(-8iJF\Delta-JF\gamma_{1}-2\gamma_{2}JF\right)(dt)^{3}$
$4iJ^{2}F(dt)^{3}$
The NESS requires $i[H_{MF,l},\rho_{l}^{SS}]=\mathcal{D}_{l}[\rho_{l}^{SS}]$, where the left-hand-side is the net coherent change in particle population at site $l$ and the right-hand-side is the incoherent loss of particles from site $l$ into the external environment (the bath). Thus it is a balance between the effects of $\Psi_{l-1}$ and $\Psi_{l+1}$ which represent the particles hopping forward from l1l-1 to $l$ and backwards from $l+1$ to $l$, respectively.
The Lindbladian dynamics of operator $a_{r}$ reads
[…]
$=\text{Tr}\left(\left(i[\hat{H},a_{r}]+\sum_{k}\left(L_{k}^{\dagger}a_{r}L_{k}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},a_{r}\}\right)\right)\hat{\rho}\right)=\text{Tr}(\mathcal{L}^{\dagger}(a_{r})\rho)=\left\langle i[H,a_{r}]+\mathcal{D}^{{\dagger}}(a_{r})\right\rangle$
$=\left\langle i\left(-\Delta\hat{a}_{l}-U\hat{n}_{l}\hat{a}_{l}\right)-iJ(-a_{r+1}-a_{r-1})-iF\delta_{1,r}-\frac{\gamma_{l}}{2}\hat{a}_{l}\right\rangle,$
where we use the cyclic property of the trace $\text{Tr}(XYZ)=\text{Tr}(ZXY)=\text{Tr}(YZX)$ and the following results
$i[\hat{H}_{\text{diag}},a_{r}]=i\left[\sum_{k}\left(\Delta n_{k}+\frac{U}{2}n_{k}(n_{k}-1)\right),a_{r}\right]=i\left[\Delta n_{r}+\frac{U}{2}(n_{r}^{2}-n_{r}),a_{r}\right]$ (22)
$=i\left(\Delta[\hat{n}_{l},\hat{a}_{l}]+\frac{U}{2}[\hat{n}_{l}^{2},\hat{a}_{l}]-\frac{U}{2}[\hat{n}_{l},\hat{a}_{l}]\right)=i\left(\Delta(-\hat{a}_{l})+\frac{U}{2}(-2\hat{n}_{l}\hat{a}_{l}-\hat{a}_{l})-\frac{U}{2}(-\hat{a}_{l})\right)$
$=i\left(-\Delta\hat{a}_{l}-U\hat{n}_{l}\hat{a}_{l}-\frac{U}{2}\hat{a}_{l}+\frac{U}{2}\hat{a}_{l}\right)=i\left(-\Delta\hat{a}_{l}-U\hat{n}_{l}\hat{a}_{l}\right),$
$i[\hat{H}_{\text{hop}},a_{r}]=i\left[-J(\hat{a}^{\dagger}_{r+1}\hat{a}_{r}+\hat{a}_{r}^{\dagger}\hat{a}_{r+1})-J(\hat{a}^{\dagger}_{r-1}\hat{a}_{r}+\hat{a}_{r}^{\dagger}\hat{a}_{r-1}),\hat{a}_{r}\right]$
$=-iJ(-a_{r+1}-a_{r-1}),$
$i[\hat{H}_{\text{drive}},\hat{a}_{l}]=i\left[F(\hat{a}^{\dagger}_{1}+\hat{a}_{1}),\hat{a}_{l}\right]\delta_{l,1}=-iF\delta_{1,r},$
$\mathcal{D}^{{\dagger}}(a_{r})=\gamma_{l}\hat{a}_{l}^{\dagger}\hat{a}_{l}\hat{a}_{l}-\frac{\gamma_{l}}{2}\{\hat{a}_{l}^{\dagger}\hat{a}_{l},\hat{a}_{l}\}=\gamma_{l}\hat{n}_{l}\hat{a}_{l}-\frac{\gamma_{l}}{2}(\hat{n}_{l}\hat{a}_{l}+\hat{a}_{l}\hat{n}_{l})$
$=\gamma_{l}\hat{n}_{l}\hat{a}_{l}-\frac{\gamma_{l}}{2}(\hat{n}_{l}\hat{a}_{l}+(\hat{n}_{l}+1)\hat{a}_{l})=\gamma_{l}\hat{n}_{l}\hat{a}_{l}-\frac{\gamma_{l}}{2}(2\hat{n}_{l}\hat{a}_{l}+\hat{a}_{l})=-\frac{\gamma_{l}}{2}\hat{a}_{l}.$
Then for NESS $\frac{d}{dt}\langle a_{r}\rangle_{ss}=0$ we can further obtain
$\langle\hat{a}_{l}\rangle\left(-i\Delta-\frac{\gamma_{l}}{2}\right)-iU\langle\hat{n}_{l}\hat{a}_{l}\rangle+iJ(\langle\hat{a}_{l-1}\rangle+\langle\hat{a}_{l+1}\rangle)=iF\delta_{l,1}$ (23)
or equivalently,
$\left(-i\Delta-\frac{\gamma_{l}}{2}\right)\Psi_{l}^{SS}-iU\langle\hat{n}_{l}\hat{a}_{l}\rangle_{ss}+iJ(\Psi_{l-1}^{SS}+\Psi_{l+1}^{SS})=iF\delta_{l,1}.$ (24)
where $\langle\hat{n}_{l}\hat{a}_{l}\rangle_{SS}$ is itself a function of $\Psi_{l\pm 1}^{SS}$. Thus for the drive, bulk, and drain sites, the NESS equations are
$-\left(i\Delta+\frac{\gamma_{1}}{2}\right)\Psi_{1}^{SS}-iU\langle\hat{n}_{1}\hat{a}_{1}\rangle_{SS}+iJ\Psi_{2}^{SS}=iF,$ (25)
$-\left(i\Delta+\frac{\gamma_{2}}{2}\right)\Psi_{2}^{SS}-iU\langle\hat{n}_{2}\hat{a}_{2}\rangle_{SS}+iJ(\Psi_{1}^{SS}+\Psi_{3}^{SS})=0,$
$-\left(i\Delta+\frac{\gamma_{3}}{2}\right)\Psi_{3}^{SS}-iU\langle\hat{n}_{3}\hat{a}_{3}\rangle_{SS}+iJ\Psi_{2}^{SS}=0$
The NESS is a highly excited state with $\langle n_{r}\rangle\sim 1$, and satisfies, $\frac{d\Psi_{r}^{ss}}{dt}=\text{Tr}\left[a_{r}\mathcal{L}_{r}(\rho_{r}^{ss})\right]=0$.
The analytical solutions are available only for $U=0$ (quadratic models), which read
$\Psi_{2}^{SS}=\frac{JF\Lambda_{3}}{\Lambda_{1}\Lambda_{2}\Lambda_{3}+J^{2}\Lambda_{3}+J^{2}\Lambda_{1}},$ (26)
$\Psi_{1}^{SS}=\frac{iF}{\Lambda_{1}}-\frac{iJ}{\Lambda_{1}}\Psi_{2}^{SS},$
$\Psi_{3}^{SS}=\frac{-iJ}{\Lambda_{3}}\Psi_{2}^{SS},$
where $\Lambda_{l}=-(i\Delta+\frac{\gamma_{l}}{2})$. The system becomes a driven-dissipative Harmonic oscillator when $U=0$, in which case $\hat{\rho}_{l}^{SS}=|\Psi_{l}^{SS}\rangle\langle\Psi_{l}^{SS}|$ is a coherent state and the local nonlinear correlation $\langle\hat{n}_{l}\hat{a}_{l}\rangle_{SS}=\langle\hat{a}_{l}^{\dagger}\hat{a}_{l}\hat{a}_{l}\rangle_{SS}=\langle\hat{a}_{l}^{\dagger}\rangle_{SS}\langle\hat{a}_{l}\rangle_{SS}\langle\hat{a}_{l}\rangle_{SS}=(\Psi_{l}^{SS})^{*}(\Psi_{l}^{SS})^{2}$ can be calculated as
$\langle\hat{n}_{l}\hat{a}_{l}\rangle$ $=\text{Tr}[\hat{\rho}_{l}\hat{a}_{l}^{\dagger}\hat{a}_{l}\hat{a}_{l}]=\sum_{n,m}\langle m|(\hat{a}_{l}^{\dagger}\hat{a}_{l}\hat{a}_{l})|n\rangle\langle n|\hat{\rho}_{l}|m\rangle$ (27)
$=\sum_{n,m}c_{n,m}\langle m|(\hat{a}_{l}^{\dagger}\hat{a}_{l}\hat{a}_{l})|n\rangle=\sum_{n=1}^{N_{\max}}c_{n-1,n}\cdot(n-1)\sqrt{n}$
where we use $\hat{a}_{l}^{\dagger}\hat{a}_{l}\hat{a}_{l}|n\rangle=(n-1)\sqrt{n}|n-1\rangle$. Note that the complex number $\Psi_{l}^{SS}=\langle\hat{a}_{l}\rangle_{SS}$ describes the amplitude and phase of boson field, with $\langle\hat{n}_{l}\rangle\approx|\Psi_{l}^{SS}|^{2}$ at small $U$ (as shown in the TWA discussed below), and $\arg(\Psi_{l}^{SS})$ determines the coherent coupling between sites.
IV Numerical Solution for NESS in the Non-Linear Regime ($U\neq 0$)
For nonzero on-site Kerr interaction $U\neq 0$, and the nonlinear correlation is $\hat{\rho}_{l}^{SS}=\sum_{n,m=0}^{N_{\text{max}}}c_{n,m}|n\rangle_{l}\langle m|_{l}$ and $\langle\hat{n}_{l}\hat{a}_{l}\rangle=\text{Tr}[\hat{\rho}_{l}\hat{a}_{l}^{\dagger}\hat{a}_{l}\hat{a}_{l}]=\sum_{n}c_{n,n-1}\langle n-1|(\hat{a}_{l}^{\dagger}\hat{a}_{l}\hat{a}_{l})|n\rangle=\sum_{n}c_{n,n-1}\cdot\sqrt{n}\cdot(n-1)$ where $c_{n,n-1}=\langle n|\rho_{r}|n-1\rangle$ and we use $\hat{a}_{l}^{\dagger}\hat{a}_{l}\hat{a}_{l}|n\rangle=\sqrt{n}(n-1)|n-1\rangle$.
The chaos is born from strong particle interactions that scramble the quantum state, and amplitudes of NESS can be solved by the above-introduced Gutzwiller mean-field method. The steady-state mean-field amplitudes $\Psi_{l}^{\text{SS}}$ of the driven-dissipative Bose-Hubbard chain are determined self-consistently using the Gutzwiller mean-field (GMF) approximation.
The effective local Hamiltonian $\hat{H}_{l}^{\text{eff}}$ for site $l$ depends on the mean-field amplitudes $\Psi_{l\pm 1}$ of its neighbors,
$\hat{H}_{l}^{\text{eff}}(\Psi)=\hat{H}_{l}^{\text{loc}}+\hat{H}_{l}^{\text{drive}}+\hat{H}_{l}^{\text{MF}}(\Psi),$ (28)
where the hopping term is linearized to the mean-field drive term $\hat{H}_{l}^{\text{MF}}(\Psi)=-J\left[(\Psi_{l-1}+\Psi_{l+1})\hat{a}_{l}^{\dagger}+(\Psi_{l-1}^{*}+\Psi_{l+1}^{*})\hat{a}_{l}\right]$, $H_{l}^{{\rm drive}}=F\delta_{l,1}(a^{{\dagger}}_{l}+a_{l})$. Here $\hat{H}_{l}^{\text{eff}}(\Psi)$ is a $N\times N$ matrix with $N=N_{max}+1$ and $N_{max}$ is the truncated number of photons. For our three-site model, $N_{max}=3$, $N=4$, thus the single-site density $\hat{\rho}_{l}$ is $N\times N$ matrix, and the dimension of total Hilbert space ($\rho_{tot}$) is $N^{N_{max}}=64$. The local steady-state density matrix $\hat{\rho}_{l}^{\text{SS}}$ is the solution to the time-independent master equation,
$\mathcal{L}_{l}(\hat{\rho}_{l}^{\text{SS}})=-i[\hat{H}_{l}^{\text{eff}}(\Psi^{(k)}),\hat{\rho}_{l}^{\text{SS}}]+\gamma_{l}\left(\hat{a}_{l}\hat{\rho}_{l}^{\text{SS}}\hat{a}_{l}^{\dagger}-\frac{1}{2}\{\hat{a}_{l}^{\dagger}\hat{a}_{l},\hat{\rho}_{l}^{\text{SS}}\}\right)=0.$ (29)
Using Roth’s Lemma ($\text{vec}(ABC)=(C^{T}\otimes A)\text{vec}(B)$), the Liouvillian reads
$\mathbf{L}\hat{\rho}_{l}^{\text{SS}}=-i(I\otimes\hat{H}_{l}^{\text{eff}}(\Psi^{(k)})-[\hat{H}_{l}^{\text{eff}}(\Psi^{(k)})]^{T}\otimes I)+\gamma_{l}(\hat{a}_{l}^{*}\otimes\hat{a}_{l}-\frac{1}{2}(I\otimes\hat{a}_{l}^{\dagger}\hat{a}_{l}+(\hat{a}_{l}^{\dagger}\hat{a}_{l})^{T}\otimes I)).$ (30)
The self-consistency condition is enforced by requiring the output amplitude $\Psi_{l}^{\text{new}}$ derived from the solution must equal the input amplitude $\Psi_{l}^{\text{old}}$ (appears in Eqs.28,29), $\Psi_{l}^{\text{new}}=\langle a_{l}\rangle_{SS}=\text{Tr}[\hat{a}_{l}\hat{\rho}_{l}^{\text{SS}}(\Psi_{l}^{\text{old}})]$, and the nonlinear effect from $U$ is contained within the $\langle n_{r}\hat{a}_{l}\rangle=\text{Tr}\left[n_{r}\hat{a}_{l}\hat{\rho}_{l}^{\text{SS}}\right]$ as well as $\hat{\rho}_{l}^{\text{SS}}$ solved from Eq.29. The algorithm iteratively using the Picard method is shown in ALGORITHM 1, where we repeatedly applying the nonlinear mapping function $\mathcal{F}(\Psi)$: $\Psi_{\text{new}}=\mathcal{F}(\Psi_{\text{old}})$, until convergence in NESS. Instead of solving the analytical expression of $\langle\hat{n}_{l}\hat{a}_{l}\rangle$, the effect of $\langle\hat{n}_{l}\hat{a}_{l}\rangle$ is captured by the matrix structure of $\Psi_{l\pm 1}^{\text{SS}}$. The Picard algorithm solve the $\hat{\rho}_{l}^{\text{SS}}$ under the effects of $U$ and $\Psi_{l\pm 1}^{\text{SS}}$. Here the coefficients of operator $\mathbf{L}$ depend on the solution $\hat{\rho}_{l\pm 1}$ (or $\Psi_{l\pm 1}$), thus $\mathbf{L}(\hat{\rho}_{l\pm 1})\hat{\rho}_{l}^{\text{SS}}=0$ is linear equation once variable of $\mathbf{L}$ is fixed by the previous iteration. Picard iteration is required to linearize the problem temporarily by fixing a guess for the neighbor fields’ amplitude.
Unlike simple mean-field theories that approximate $\langle\hat{n}\hat{a}\rangle\approx\langle\hat{n}\rangle\langle\hat{a}\rangle$, GMF handles the non-linearity ($U\neq 0$) locally by solving $\mathcal{L}_{l}(\hat{\rho}_{l})=0$. The non-linearity is moved from the operators (Liouvillian) to the iteration loop, i.e., the nonlinear mapping $\Psi_{\text{old}}\to\Psi_{\text{new}}$ as $\hat{\rho}_{l}$ depends on $\Psi_{\text{old}}$ in a complex and non-proportional way (e.g., through the photon blockade).
Algorithm 1 Self-Consistent GMF NESS Solver (Picard Iteration)
1:  Input: Physical parameters $(J,U,\Delta,F,\{\gamma_{l}\})$; Convergence tolerance $\epsilon$; Max Iterations $M$.
2:  Initialize: Initial guess $\Psi_{\text{old}}^{(0)}=\{\Psi_{1},\Psi_{2},\Psi_{3}\}$; Iteration counter $k\leftarrow 0$.
3:  while $k<M$ do
4:   Compute New Amplitudes $\Psi_{\text{new}}\leftarrow\mathcal{F}(\Psi_{\text{old}})$:
5:   for each site $l=1$ to $L$ do
6:    $\hat{H}_{l}^{\text{eff}}\leftarrow$ Compute $\hat{H}_{l}^{\text{eff}}$ using $\Psi_{l\pm 1}^{\text{old}}$.
7:    $\mathbf{L}_{l}\leftarrow$ Construct the vectorized Liouvillian matrix for $\hat{H}_{l}^{\text{eff}}$.
8:    $\vec{\rho}_{l}^{\text{SS}}\leftarrow$ Solve $\mathbf{L}_{l}\vec{\rho}_{l}=0$ subject to $\text{Tr}(\hat{\rho}_{l})=1$ (which becomes $\mathbf{L}^{\prime}_{l}\vec{\rho}_{l}\neq 0$ under trace trick).
9:    $\Psi_{l}^{\text{new}}\leftarrow$ Compute $\text{Tr}[\hat{a}_{l}\hat{\rho}_{l}^{\text{SS}}]$.
10:   end for
[…]
12:   if $R<\epsilon$ then
13:    break (Solution $\Psi_{\text{SS}}=\Psi_{\text{new}}$ Found)
14:   end if
15:   Update: $\Psi_{\text{old}}\leftarrow\Psi_{\text{new}}$; $k\leftarrow k+1$.
16:  end while
The trace conservation requires $\frac{d}{dt}\text{Tr}(\hat{\rho})=\text{Tr}\left(\frac{d\hat{\rho}}{dt}\right)=\langle\langle\mathbb{1}|\frac{d}{dt}\rho\rangle\rangle=\langle\langle\mathbb{1}|\mathbf{L}|\rho\rangle\rangle=0$, where we use the Hilbert-Schmidt inner product of the form $\text{Tr}(\hat{A})=\text{Tr}(\hat{\mathbb{1}}^{\dagger}\hat{A})\equiv\langle\langle\mathbb{1}|A\rangle\rangle$ with the vectorized density $|A\rangle\rangle$ (according to Roth’s Lemma). Thus $\langle\langle\mathbb{1}|\mathbf{L}=\mathbf{0}^{T}$, $\det(\mathbf{L})=\prod\lambda_{i}=0$, and consequently $\mathbf{L}$ is a singular matrix. Meanwhile the singularity of $\mathbf{L}$ guarantees that $\mathbf{L}|\rho_{ss}\rangle\rangle=0$ has nontrivial solution (nonzero steady state). Further, since $\frac{d}{dt}\text{Tr}(\hat{\rho})=\langle\langle\mathbb{1}|\frac{d}{dt}|\rho\rangle\rangle=\langle\langle\mathbb{1}|(\mathbf{L}|\rho\rangle\rangle)=(\langle\langle\mathbb{1}|\mathbf{L})|\rho\rangle\rangle=0$, where the vector $\langle\langle\mathbb{1}|$ is equivalents to $|\rho\rangle\rangle^{T}$ but replace the diagonal element $\rho_{ii}$ by 1 and replace the off-diagonal element $\rho_{ij}$ by 0. Then we have $\langle\langle\mathbb{1}|\mathbf{L}=\vec{0}^{T}$, which means that for each column of $\mathbf{L}$, sum of elements corresponding to the positions of nonzero elements of $\langle\langle\mathbb{1}|$ must be zero, such that the sum of the corresponding rows of $\mathbf{L}$ is zero (and thus those rows are linearly dependent):
$L_{1,j}+L_{N+2,j}+L_{2N+3,j}+\cdots+L_{N(N-1)+N,j}=0,$ (31)
with $N$ the dimension of matrix $\rho$. This is equivalents to the trace conservation condition
$\frac{d}{dt}\text{Tr}(\rho)=\sum_{i}\frac{d\rho_{ii}}{dt}=\sum_{i}\frac{d|\rho\rangle\rangle_{i}}{dt}=\sum_{i}(\mathbf{L}|\rho\rangle\rangle)_{i}=\sum_{i}\sum_{j=1}^{N^{2}}L_{i,j}|\rho\rangle\rangle_{j}=\sum_{j=1}^{N^{2}}|\rho\rangle\rangle_{j}\left(\sum_{i}L_{i,j}\right)=0,$ (32)
where here $i=1,N+2,2N+3,\dots,N(N-1)+N$ corresponds to the position of the diagonal elements. Multiple with $|\rho\rangle\rangle$ in Eq.31, we can again obtain $\frac{d\rho_{11}}{dt}+\dots\frac{d\rho_{NN}}{dt}=0$. For singular matrix $\mathbf{L}$, $\mathbf{L}\vec{\rho}=\mathbf{0}$ describes a homogeneous linear system, thus introducing the condition $\text{Tr}(\hat{\rho})=1$ is necessary to find the sole solution. This can be realized by replacing the first row of $\mathbf{L}$ by the vector satisfying $\langle\langle\mathbb{1}|\rho\rangle\rangle=1\cdot\rho_{11}+0\cdot\rho_{12}+\dots+1\cdot\rho_{22}+\dots=\sum_{i}\rho_{ii}={\rm Tr}\rho=1$ (guarantees the normalization condition), while for other rows the product with $|\rho\rangle\rangle$ remains zero. The modification of first row of $\mathbf{L}$ realize
$\frac{d|\rho\rangle\rangle}{dt}=\mathbf{L}|\rho\rangle\rangle=\begin{pmatrix}\frac{d\rho_{11}}{dt}\\ \frac{d\rho_{12}}{dt}\\ \vdots\\ \frac{d\rho_{NN}}{dt}\end{pmatrix}\rightarrow\mathbf{L}^{\prime}|\rho\rangle\rangle=\begin{pmatrix}\sum_{i}\rho_{ii}\\ \frac{d\rho_{12}}{dt}\\ \vdots\\ \frac{d\rho_{NN}}{dt}\end{pmatrix}=\begin{pmatrix}1\\ \frac{d\rho_{12}}{dt}\\ \vdots\\ \frac{d\rho_{NN}}{dt}\end{pmatrix}=\begin{pmatrix}1\\ 0\\ \vdots\\ 0\end{pmatrix}$. (33)
For full rank invertible (nonsinglular) matrix $\mathbf{L}^{\prime}$, the non-homogeneous linear system $\mathbf{L}^{\prime}|\rho\rangle\rangle$ has sole solution: $\mathbf{L}^{\prime}|\rho\rangle\rangle=0$ has only one solution $|\rho\rangle\rangle=0$, $\mathbf{L}^{\prime}|\rho\rangle\rangle=b\neq 0$ has only one solution $|\rho\rangle\rangle=\mathbf{L}^{{}^{\prime}-1}b$. For singlular matrix $\mathbf{L}$, the homogeneous linear system $\mathbf{L}|\rho\rangle\rangle$ has infinite solutions: $\mathbf{L}|\rho\rangle\rangle=0$ has a solution $|\rho\rangle\rangle=0$, and infinite other nonzero solutions $c|\rho\rangle\rangle_{ss}$ ($c$ is arbitary constant), and all these solutions form the so-called null space or kernel. $\mathbf{L}|\rho\rangle\rangle=b\neq 0$ has infinite nonzero solutions if $\vec{b}$ lies in the column space of $\mathbf{L}$. Thus for homogeneous linear system $\mathbf{L}|\rho\rangle\rangle$, if solutions $|\rho_{1}\rangle\rangle=c|\rho_{1}\rangle\rangle$, then ${\rm Tr}\rho_{1}=c{\rm Tr}\rho_{2}$. Note that $\mathbf{L}$ in our system (or other ergodic open quantum system) always has rank $(N^{2}-1)$ which means the dimension of nullspace is 1 (this dimension will be larger than 1 when there are dark states or symmetry-protected/decoherence-free subspaces), corresponding to the number of unique steady state. Further, for singlular matrix $\mathbf{L}$, the nonzero solution for $\mathbf{L}|\rho\rangle\rangle$ could have arbitary trace ${\rm Tr}\rho_{ss}$ as long as $||\rho_{ss}\rangle\rangle|^{2}=1$. ${\rm Tr}\rho_{ss}$ could even be zero when there are dark states or symmetry-protected/decoherence-free subspaces (like the case $J=0$). For $\mathbf{L}$ with rank $(N^{2}-1)$, there exists one vector $\langle\langle\mathbb{1}|$ satisfying $\langle\langle\mathbb{1}|\mathbf{L}=\vec{0}^{T}$, for special $\mathbf{L}$ with rank $(N^{2}-k)$, there are $k$ linearly independent steady states, i.e., expect vector $\langle\langle\mathbb{1}|$, there are additionally k1k-1 vector $v$ satisfying $v\mathbf{L}=\vec{0}^{T}$, and the final state depends on initial conditions. In this case, there exist $k$ linearly independent left eigenvectors with eigenvalue 0, representing $k$ independent conservation laws (one being the total trace, others being e.g. population in decoupled subspaces).
For the special $\mathbf{L}$ with rank $(N^{2}-k)$ ($k>1$), there is a $k$-dimensional nullspace spanned by $d$ linearly independent basis vectors. We consider the conservation law of a observable $O$
$\frac{d}{dt}\langle\hat{O}\rangle=\text{Tr}\left(\hat{O}\frac{d\hat{\rho}}{dt}\right)=\text{Tr}\left(\hat{O}\mathcal{L}(\hat{\rho})\right)=\text{Tr}(\hat{O}^{\dagger}\mathcal{L}(\hat{\rho}))=\langle\langle O|\mathcal{L}(\hat{\rho})\rangle\rangle=\langle\langle O|\left(\mathbf{L}|\rho\rangle\rangle\right)=\left(\langle\langle O|\mathbf{L}\right)|\rho\rangle\rangle=\vec{v}\mathbf{L}|\rho\rangle\rangle,$ (34)
[…]
While the mathematical procedure ensures $\text{Tr}(\hat{\rho})=1$, numerical floating-point errors may introduce a negligible deviation (on the order of machine epsilon, $\sim 10^{-16}$). It is safe and sometimes recommended to explicitly re-normalize the density matrix after reshaping it from the vector form:
$\hat{\rho}\leftarrow\frac{\hat{\rho}}{\text{Tr}(\hat{\rho})}$ (35)
This step acts as a safeguard against accumulated numerical errors but is strictly redundant in exact arithmetic. Note that as long as $F\neq 0$ or $\Delta\neq 0$, $\rho$ is a complex Hermitian matrix and positive semi-definite with all eigenvalues be nonnegative.
In GMF approximation, the total density matrix reads $\hat{\rho}_{\text{tot}}^{\text{GMF}}=\hat{\rho}_{1}\otimes\hat{\rho}_{2}\otimes\dots\otimes\hat{\rho}_{L}$. The interaction between site $l$ and its neighbor $l+1$ is replaced by an effective field $J\hat{a}_{l}^{\dagger}\hat{a}_{l+1}\approx-J\hat{a}_{l}^{\dagger}\langle\hat{a}_{l+1}\rangle$, where $\langle\hat{a}_{l+1}\rangle=\text{Tr}(\hat{a}_{l+1}\hat{\rho}_{l+1})$ depends on the density matrix $\hat{\rho}_{l+1}$ solved self-consistency loop $\mathbf{L}(\hat{\rho})\cdot\hat{\rho}=0$. Thus the hopping term coefficient $J\Psi_{l+1}=J{\rm Tr}[\hat{a}_{l+1}\hat{\rho}_{l+1}]$ depends on $\hat{\rho}_{l+1}$. Even though $U,J,F,\gamma$ are constant, the coefficients of the hopping operator in the matrix $\mathbf{L}$ depend on the solution $\hat{\rho}$ of the adjacent sites, i.e., to solve for $\rho_{l}$, $\Psi_{l+1}$ comes from $\rho_{l+1}$ is needed, and vise-vase. Picard iteration breaks this circle by fixing $\rho$ from the previous step to calculate the hopping coefficients for the current step.
For a comparasion, we also apply the exact solver use the global basis space with size $N^{N_{max}}$, where the NESS solver constructs the Liouvillian in the full $(N^{N_{max}})^{2}$ space. The interaction term $\frac{U}{2}\hat{n}(\hat{n}-1)$ is represented as a static $(N^{N_{max}})\times(N^{N_{max}})$ matrix operator. The hopping term $J\hat{a}_{l}^{\dagger}\hat{a}_{l+1}$ is also a static $(N^{N_{max}})\times(N^{N_{max}})$ matrix operator. In the master equation
$\frac{d\hat{\rho}_{\text{tot}}}{dt}=-i[\hat{H}_{\text{tot}},\hat{\rho}_{\text{tot}}]+\sum\mathcal{D}[\hat{L}](\hat{\rho}_{\text{tot}}),$ (36)
with constant coefficients $U,J,F,\gamma$ that do not depend on the value of $\hat{\rho}_{\text{tot}}$. Thus, the equation is a standard linear system $\mathbf{L}_{\text{tot}}|\rho_{\text{tot}}\rangle\rangle=0$. Despite the non-linearity $U$, the density matrix equation itself is linear in the state space. Therefore the NESS can be solved without iteration.
V OTOC
In the Gutzwiller mean-field framework linearized around the NESS, this OTOC is related to the squared magnitude of the retarded Green’s function (propagator) governing the fluctuations $\delta\hat{a}$:
$D_{1,L}(\tau)\approx|G^{R}_{L,1}(\tau)|^{2}=\left|\left(e^{\mathbf{L}_{\text{chain}}\tau}\right)_{L,1}\right|^{2},$ (37)
where $\mathbf{L}_{\text{chain}}$ is the collective effective Liouvillian for the coupled chain fluctuations. Exponential growth of $D_{1,L}(\tau)$ indicates dynamical instability and the onset of chaos.
Next we consider the out-of-time-ordered correlator (OTOC)
$\mathbf{D}_{ij}(\tau)=\text{Tr}\left[\hat{\rho}_{SS}[a_{i}(\tau),a^{{\dagger}}_{j}(0)]^{2}\right]=\langle[\hat{a}_{i}(\tau),\hat{a}_{j}^{\dagger}(0)]^{\dagger}[\hat{a}_{i}(\tau),\hat{a}_{j}^{\dagger}(0)]\rangle_{SS}$ (38)
$=\langle\hat{a}_{j}(0)\hat{a}_{i}^{{\dagger}}(\tau)\hat{a}_{i}(\tau)\hat{a}_{j}^{\dagger}(0)-\hat{a}_{j}(0)\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{j}^{\dagger}(0)\hat{a}_{i}(\tau)-\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{j}(0)\hat{a}_{i}(\tau)\hat{a}_{j}^{\dagger}(0)+\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{j}(0)\hat{a}_{j}^{\dagger}(0)\hat{a}_{i}(\tau)\rangle$
$=\frac{2\langle\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{i}(\tau)\rangle\langle\hat{a}_{j}^{\dagger}(0)\hat{a}_{j}(0)\rangle-2\text{Re}(\langle\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{j}(0)\hat{a}_{i}(\tau)\hat{a}_{j}^{\dagger}(0)\rangle)}{\langle\hat{a}_{i}^{\dagger}(\tau)\hat{a}_{i}(\tau)\rangle\langle\hat{a}_{j}^{\dagger}(0)\hat{a}_{j}(0)\rangle}$
$=2\left(1-\text{Re}\left[\frac{\langle a_{i}^{\dagger}(\tau)a_{j}a_{i}(\tau)a^{{\dagger}}_{j}\rangle_{SS}}{\langle a_{i}^{\dagger}a_{i}\rangle_{SS}\langle a_{j}a^{{\dagger}}_{j}\rangle_{SS}}\right]\right)$
[…]
In the weak drive/linear limit, commutators decay or remain constant. $D_{1,L}(\tau\to\infty)\approx 0$. While in the strong coupling regime ($U\gg\gamma$), the non-linearity scrambles local information globally. The OTOC exhibits characteristic exponential growth at early times, $D(\tau)\sim e^{\lambda_{L}\tau}$ (where $\lambda_{L}$ is the quantum Lyapunov exponent), before saturating to a finite value $O(1)$ determined by the local Hilbert space dimension. This saturation indicates that information about the drive at site 1 has been effectively "forgotten" or scrambled across the entire many-body Hilbert space of the chain.
Fig.2 show the OTOC as a function of time in unit of inverse hopping amplitude ($1/J$; such that the value $t=1.0$ corresponds to the characteristic time required for a particle to tunnel between neighboring sites). for site 3 (drain site) is of the strong dissipative and quasilinear regime, Although quantum information propagates to this site as indicated by the initial rise, it is quickly lost to the environment due to strong coupling with the bath. The vanishing saturation value is consistent with a strong dissipation or quasilinear regime, where information is not scrambled locally but rather "flushed out" before it can accumulate.
While site 1 and site 2 are of the chaotic regime. For $t<0.5$, there is exponential growth as a hallmark of the Lyapunov regime, indicating that the nonlinear interaction $U$ is actively scrambling local information. The peak for site 2 at $t\approx 0.5$ is consistent as it takes roughly half a tunneling period for the excitation to propagate from the drive (site 1) to the bulk (site 2). The OTOC does not saturate smoothly but exhibits large-amplitude oscillations. This reflects strong coherent interactions which is typical for finite-size systems. The wave packet reflects off the boundaries, causing recurrences. Unlike the Drain site, the OTOC for the Drive and Bulk sites does not decay to zero. Instead, it fluctuates around a finite value ($O(1)$ to $O(3)$) at long times ($t=5$). Saturation to a finite value at long time signaling significant dephasing. The high peak in the bulk site suggests strong parametric amplification of fluctuations due to nonlinearity before dissipation takes over. The fact that the signal remains non-zero suggests the system is in the chaotic regime, where operators $\hat{a}(t)$ and $\hat{a}^{\dagger}(0)$ fail to commute over long periods, signifying sustained scrambling despite the presence of dissipation. The competition between unitary scrambling (driven by nonlinearity $U$) and non-unitary dissipation (driven by $\gamma$).
The phase OTOC evolution reveals a distinct separation of timescales. We observe an initial Lyapunov-like exponential growth driven by the Kerr nonlinearity, followed by bounded oscillations characteristic of finite-size coherent dynamics. Crucially, the bulk and drive sites maintain non-zero correlations at late times, indicative of the chaotic regime, whereas the drain site exhibits rapid decay due to strong coupling with the environment.
[…]
Figure 2: (Left) Time evolution of the local Phase OTOC $D_{3,3}(t)$ calculated on NESS background using the Picard iteration self-consistent solver. The rapid initial growth and subsequent saturation at a non-zero value indicate significant information scrambling and the onset of local quantum chaos induced by the Kerr nonlinearity. (Right) The OTOC solved by exact solver.
Figure 3: The population base on TWA for 100 trajectories (left) and 500 trajectories (right). The blue, orange, and green correspond to the site 1, site 2, and site 3, respectively. At long time, $\langle n_{1}\rangle\approx\langle n_{3}\rangle\approx\frac{3}{2}\langle n_{2}\rangle$.
VI Effective Liouvillian $\mathbf{L}_{eff}$
[…]
$\frac{U}{2}(\Psi_{SS}^{*}+\delta\hat{a}^{\dagger})^{2}(\Psi_{SS}+\delta\hat{a})^{2}\approx\frac{U}{2}\left[(\Psi_{SS}^{*})^{2}+2\Psi_{SS}^{*}\delta\hat{a}^{\dagger}+(\delta\hat{a}^{\dagger})^{2}\right]\left[\Psi_{SS}^{2}+2\Psi_{SS}\delta\hat{a}+(\delta\hat{a})^{2}\right].$ (39)
The constant terms only contributes to the energy offset and does not affect dynamics. The linear terms ($\propto\delta\hat{a}$) vanish identically because $\Psi_{SS}$ is defined as the steady state where the net force should be zero (corresponding to the saddle point of the action). If the sum of all linear driving forces is finite, $\Psi_{SS}$ would change over time, violates the condition of NESS.
Only the quadratic leading terms determine the linearized dynamics (small oscillations) and determine stability. Higher order terms describing the interactions between the small fluctuations are also ignored. Retaining terms up to quadratic order in fluctuations $\delta\hat{a}$ which approximates the potential energy surface around the steady state as a harmonic oscillator, the fluctuation dynamics reads
$\hat{H}_{\text{fluct}}\approx(\Delta+2U|\Psi_{SS}|^{2})\delta\hat{a}^{\dagger}\delta\hat{a}+\frac{U}{2}\Psi_{SS}^{2}(\delta\hat{a}^{\dagger})^{2}+\frac{U}{2}(\Psi_{SS}^{*})^{2}(\delta\hat{a})^{2}.$ (40)
The term $2U|\Psi_{SS}|^{2}$ represents the Hartree-Fock energy shift, while the terms proportional to $\Psi_{SS}^{2}$ represent parametric driving terms[3, 4, 5, 6] that allow the creation or annihilation of pairs of fluctuations from vacuum driven by the coherent field. Higher orders terms representing the interactions between fluctuations are neglected to obtain a solvable linear system (Gaussian approximation) for stability analysis.
The time evolution is governed by the Heisenberg-Langevin equation, $\frac{d}{dt}\delta\hat{a}=-i[\delta\hat{a},\hat{H}_{\text{fluct}}]-\frac{\gamma}{2}\delta\hat{a}+\hat{\xi}(t)$, $\frac{d}{dt}\delta\hat{a}^{{\dagger}}=-i[\delta\hat{a}^{{\dagger}},\hat{H}_{\text{fluct}}]-\frac{\gamma}{2}\delta\hat{a}^{{\dagger}}+\hat{\xi}(t)$. $\hat{\xi}(t)$ is quantum noise (Langevin noise) arising from the coupling to the environment as required by the fluctuation-dissipation theorem. Without this noise term, the canonical commutation relation $\delta\hat{a}(t),\delta\hat{a}^{\dagger}(t)]$ would decay to zero due to the damping $\frac{\gamma}{2}\delta\hat{a}$, violating quantum mechanics. The noise correlation $\langle\hat{\xi}(t)\hat{\xi}^{\dagger}(t^{\prime})\rangle=\gamma\delta(t-t^{\prime})$ continuously replenishes the quantum fluctuations, ensuring $\delta\hat{a}(t),\delta\hat{a}^{\dagger}(t)]=1$ at all times.
It is necessary to preserve the commutation relation $\hat{a}(t),\hat{a}^{\dagger}(t)]=1$ as the system decays. Calculating the commutators $\delta\hat{a},\delta\hat{a}^{\dagger}\delta\hat{a}]=\delta\hat{a}$ and $\delta\hat{a},(\delta\hat{a}^{\dagger})^{2}]=2\delta\hat{a}^{\dagger}$, we obtain the linearized equation of motion for the deterministic drift (neglecting the noise term $\hat{\xi}(t)$ for the eigenvalue stability analysis)
$\frac{d}{dt}\delta\hat{a}=-i\left((\Delta+2U|\Psi_{SS}|^{2})\delta\hat{a}+U\Psi_{SS}^{2}\delta\hat{a}^{\dagger}\right)-\frac{\gamma}{2}\delta\hat{a}$, (41)
$\frac{d}{dt}\delta\hat{a}^{{\dagger}}=i\left((\Delta+2U|\Psi_{SS}|^{2})\delta\hat{a}^{{\dagger}}+U(\Psi_{SS}^{*})^{2}\delta\hat{a}\right)-\frac{\gamma}{2}\delta\hat{a}^{{\dagger}}.$
The above linearized equation of motion describes the deterministic drift of the fluctuations where the noise term $\hat{\xi}(t)$ was averaged out. $\hat{\xi}(t)$ disappears in the matrix form because the matrix $\mathbf{L}_{eff}$ describes the drift (deterministic) part of the evolution. Stability is determined by the eigenvalues of this drift matrix. If the drift causes fluctuations to decay (stable eigenvalues), the noise maintains a finite variance. If the drift causes growth (unstable eigenvalues), the noise is amplified exponentially. Only the homogeneous part of the differential equation $\mathbf{L}_{eff}$ determine if the NESS is stable. If the eigenvalues of deterministic matrix have negative real parts, any perturbation decays. If they have positive real parts, the perturbation grows. The random noise $\hat{\xi}(t)$ acts as an inhomogeneous driving term and does not determine whether those modes are stable or unstable, and thus it is dropped when calculating the eigenvalues.
The stability of a NESS is only determined by the homogeneous part of the differential equation ($\mathbf{L}_{eff}$). If the eigenvalues of the deterministic matrix have negative real parts, any perturbation decays. If they have positive real parts, the perturbation grows. The quantum expectation value of the Langevin equation reads
$\frac{d}{dt}\langle\delta\hat{a}\rangle$ $=-i\left((\Delta+2U|\Psi_{SS}|^{2})\delta\hat{a}+U\Psi_{SS}^{2}\delta\hat{a}^{\dagger}\right)-\frac{\gamma}{2}\langle\delta\hat{a}\rangle+\langle\hat{\xi}(t)\rangle$ (42)
$=-i\left((\Delta+2U|\Psi_{SS}|^{2})\delta\hat{a}+U\Psi_{SS}^{2}\delta\hat{a}^{\dagger}\right)-\frac{\gamma}{2}\langle\delta\hat{a}\rangle$
since $\langle\hat{\xi}(t)\rangle=0$).
Different to above nonlinear equation $\frac{d}{dt}\hat{a}_{r}$, analytical solutions are available here, which can be expressed as the superposition of two exponential modes corresponding to the eigenvalues $\lambda_{\pm}$,
$\delta\hat{a}(t)=u(t)\delta\hat{a}(0)+v(t)\delta\hat{a}^{\dagger}(0)$ (43)
with
$u(t)$ $=e^{-\frac{\gamma}{2}t}\left[\cos(\Omega t)-i\frac{\Delta+2U|\Psi_{SS}|^{2}}{\Omega}\sin(\Omega t)\right]$ (44)
$v(t)$ $=e^{-\frac{\gamma}{2}t}\left[-i\frac{U\Psi_{SS}^{2}}{\Omega}\sin(\Omega t)\right],$
$\Omega=\sqrt{(\Delta+2U|\Psi_{SS}|^{2})^{2}-|U\Psi_{SS}^{2}|^{2}}$
where $|\Delta+2U|\Psi_{SS}|^{2}|$ is the effective detuning, $U\Psi_{SS}^{2}$ is the parametric gain, and $\Omega$ is the growth rate or oscillation frequency. Then for $(\Delta+2U|\Psi_{SS}|^{2})^{2}>|U\Psi_{SS}^{2}|^{2}$ (stable and oscillatory), $\Omega$ is real and $\sinh(\Omega t)\sim\frac{1}{2}e^{\Omega t}$. The solution is a damped oscillation decaying at rate $\gamma/2$. For $|U\Psi_{SS}^{2}|^{2}>(\Delta+2U|\Psi_{SS}|^{2})^{2}$ (unstable and parametric), $\Omega$ is imaginary. If this growth rate $\Omega$ is faster than the decay $\gamma/2$ (i.e., $\Omega>\gamma/2$), the fluctuation $\delta\hat{a}(t)$ will explode exponentially. The off-diagonal term $v(t)$ and $v^{*}(t)$ critically depends on the ononlinear strength $U$. In linear limit $U=0$, $U\Psi_{SS}^{2}=0$, $\Delta+2U|\Psi_{SS}|^{2}=\Omega=\Delta$, and $u(t)=e^{-\frac{\gamma}{2}t}(\cos\Delta t-i\sin\Delta t)=e^{-i\Delta t-\frac{\gamma}{2}t}$, $v(t)=0$. Thus $\delta\hat{a}(t)=\delta\hat{a}(0)e^{-i\Delta t-\frac{\gamma}{2}t}$, without squeezing.
The commutator of the deterministic part of the fluctuation reads
$\delta\hat{a}(t),\delta\hat{a}^{\dagger}(t)]=\left[u(t)\delta\hat{a}(0)+v(t)\delta\hat{a}(0)^{\dagger},u^{*}(t)\delta\hat{a}(0)^{\dagger}+v^{*}(t)\delta\hat{a}(0)\right]=|u(t)|^{2}-|v(t)|^{2}=e^{-\gamma t}.$ (45)
This shows that the system loss the quantum uncertainty of its initial state due to dissipation. For closed system with $\gamma=0$, $|u(t)|^{2}-|v(t)|^{2}=1$ corresponds to unitary evolution without loss of quantum information and preserved commutation relation. The full time evolution of the operator in the linearized regime reads
$\hat{a}(t)=\Psi_{SS}+u(t)[\hat{a}(0)-\Psi_{SS}]+v(t)[\hat{a}^{\dagger}(0)-\Psi_{SS}^{*}]+\hat{\mathcal{F}}(t)$ (46)
where the second and third terms in the right-hand-side represent the deterministic part[14] (memory of initial state). $\hat{\mathcal{F}}(t)$ is the accumulated noise integral. $u(t)$ and $v(t)$ are the entries of the matrix propagator $e^{\mathbf{L}_{\text{eff}}t}$. The coherent amplitude $\Psi_{SS}$ is the constant classical background of NESS. $u(t)$ is the normal evolution (rotation and decay), $v(t)$ is the anomalous evolution (mixing creation and annihilation operators due to the non-linearity $U$). $\hat{\mathcal{F}}(t)$ the accumulated quantum noise (integral of $\hat{\xi}(t^{\prime})$), which is necessary to keep $\hat{a}(t),\hat{a}^{\dagger}(t)]=1$ as the deterministic parts $u(t)$ and $v(t)$ decay. The total operator commutator must remain 1 for all time,
$\hat{a}(t),\hat{a}^{\dagger}(t)]=[\delta\hat{a}(t)+\hat{\mathcal{F}}(t),\delta\hat{a}^{\dagger}(t)+\hat{\mathcal{F}}^{\dagger}(t)]=[\delta\hat{a}(t),\delta\hat{a}^{\dagger}(t)]+[\hat{\mathcal{F}}(t),\hat{\mathcal{F}}^{\dagger}(t)]=e^{-\gamma t}+[\hat{\mathcal{F}}(t),\hat{\mathcal{F}}^{\dagger}(t)]=1,$ (47)
where the cross terms $\delta\hat{a}(t),\hat{\mathcal{F}}^{\dagger}(t)]$ vanish because initial operators commute with future noise operators). According to fluctuation-dissipation theorem, this implies $\hat{\mathcal{F}}(t),\hat{\mathcal{F}}^{\dagger}(t)]=1-e^{-\gamma t}$. The accumulated noise $\hat{\mathcal{F}}(t)$ exactly replenishes the commutator as it decays. Without $\hat{\mathcal{F}}(t)$, the commutator would go to zero, which is a violation of the laws of quantum mechanics.
With spinor $(\delta\hat{a},\delta\hat{a}^{\dagger})^{T}$ and the effective Liouvillian matrix $\mathbf{L}_{eff}$ (Bogoliubov-de Gennes matrix), the deterministic evolution reads
$\frac{d}{dt}\begin{pmatrix}\delta\hat{a}\\ \delta\hat{a}^{\dagger}\end{pmatrix}=\underbrace{\begin{pmatrix}-i(\Delta+2U|\Psi_{SS}|^{2})-\frac{\gamma}{2}&-iU\Psi_{SS}^{2}\\ iU(\Psi_{SS}^{*})^{2}&i(\Delta+2U|\Psi_{SS}|^{2})-\frac{\gamma}{2}\end{pmatrix}}_{\mathbf{L}_{eff}}\begin{pmatrix}\delta\hat{a}\\ \delta\hat{a}^{\dagger}\end{pmatrix}$. (48)
The complex eigenvalues $\lambda_{\pm}$ of $\mathbf{L}_{eff}$ determine the system’s stability and oscillation frequencies,
$\lambda_{\pm}=-\frac{\gamma}{2}\pm i\sqrt{(\Delta+2U|\Psi_{SS}|^{2})^{2}-|U\Psi_{SS}^{2}|^{2}}.$ (49)
The real part of the eigenvalues are
$\text{Re}(\lambda_{\pm})=\begin{cases}-\frac{\gamma}{2}&\text{if }(\Delta+2U|\Psi_{SS}|^{2})^{2}\geq|U\Psi_{SS}^{2}|^{2}\qquad\text{(stable/oscillatory regime)}\\ -\frac{\gamma}{2}\mp\sqrt{|U\Psi_{SS}^{2}|^{2}-(\Delta+2U|\Psi_{SS}|^{2})^{2}}&\text{if }(\Delta+2U|\Psi_{SS}|^{2})^{2}<|U\Psi_{SS}^{2}|^{2}\qquad\text{(prarametric Regime)}\end{cases}$ (50)
Note that $\gamma>0$, thus the NESS is stable if and only if the largest real part $\text{Re}(\lambda_{-})<0$. If the parametric gain $|U\Psi_{SS}^{2}|$ is sufficiently strong such that $\sqrt{|U\Psi_{SS}^{2}|^{2}-(\Delta+2U|\Psi_{SS}|^{2})^{2}}>\frac{\gamma}{2}$, the real part becomes positive,
$|U\Psi_{SS}^{2}|^{2}>(\Delta+2U|\Psi_{SS}|^{2})^{2}+\frac{\gamma^{2}}{4}.$ (51)
The NESS is stable if and only if all eigenvalues satisfy $\text{Re}(\lambda)<0$. If the term under the square root becomes sufficiently negative (or if the parametric gain dominates the damping). Positive $\text{Re}(\lambda)$ signaling an exponential growth of fluctuations and a transition to a dynamical instability (parametric oscillation) and chaotic regime.
Since $\det(e^{\mathbf{A}})=e^{\text{Tr}(\mathbf{A})}$ and $\text{Tr}(\mathbf{L}_{\text{eff}})=(-i\tilde{\Delta}-\frac{\gamma}{2})+(i\tilde{\Delta}-\frac{\gamma}{2})=-\gamma$, we have $\det(e^{\mathbf{L}_{\text{eff}}t})=u(t)u^{*}(t)-v(t)v^{*}(t)=|u(t)|^{2}-|v(t)|^{2}=e^{\text{Tr}(\mathbf{L}_{\text{eff}})t}=e^{-\gamma t}$. Using Cayley-Hamilton theorem, the 2-by-2 matrix $\mathbf{M}:=\mathbf{L}_{eff}+\frac{\gamma}{2}{\bf I}$ satisfy $\mathbf{M}^{2}-\text{tr}(\mathbf{M})\mathbf{M}+\det(\mathbf{M})\mathbf{I}=\mathbf{0}$, thus $\mathbf{M}^{2}=-\det(\mathbf{M})\mathbf{I}=-\Omega^{2}{\bf I}$. Note that in oscillatory stable regime, $\det(\mathbf{M})=(-i(\Delta+2U|\Psi_{SS}|^{2}))(i(\Delta+2U|\Psi_{SS}|^{2}))-(-iU\Psi_{SS}^{2})(iU(\Psi_{SS}^{*})^{2})=(\Delta+2U|\Psi_{SS}|^{2})^{2}-|U\Psi_{SS}^{2}|^{2}=\Omega^{2}\geq 0$.
$e^{\mathbf{M}t}=\sum_{n=0}^{\infty}\frac{(\mathbf{M}t)^{n}}{n!}=\mathbf{I}+\mathbf{M}t+\frac{\mathbf{M}^{2}t^{2}}{2!}+\frac{\mathbf{M}^{3}t^{3}}{3!}+\frac{{\bf M}^{4}t^{4}}{4!}+\dots$ (52)
$=\mathbf{I}+\mathbf{M}t+\frac{-\Omega^{2}{\bf I}t^{2}}{2!}+\frac{-\Omega^{2}{\bf I}\mathbf{M}t^{3}}{3!}+\frac{\Omega^{4}{\bf I}t^{4}}{4!}+\dots$
$=\sum_{k=0}^{\infty}\frac{t^{2k}}{(2k)!}(-\Omega^{2})^{k}\mathbf{I}+\mathbf{M}\sum_{k=0}^{\infty}\frac{t^{2k+1}}{(2k+1)!}(-\Omega^{2})^{k}$
$=\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}(\Omega t)^{2k}}{(2k)!}\right)\mathbf{I}+\frac{\mathbf{M}}{\Omega}\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}(\Omega t)^{2k+1}}{(2k+1)!}\right)$
$=\cos(\Omega t)\mathbf{I}+\frac{\sin(\Omega t)}{\Omega}\mathbf{M}$
$=\begin{pmatrix}\cos(\Omega t)&0\\ 0&\cos(\Omega t)\end{pmatrix}+\frac{\sin(\Omega t)}{\Omega}\begin{pmatrix}-i(\Delta+2U|\Psi_{SS}|^{2})&-iU\Psi_{SS}^{2}\\ i(U\Psi_{SS}^{2})^{*}&i(\Delta+2U|\Psi_{SS}|^{2})\end{pmatrix}$
Then we can obtain
$u(t)=e^{-\frac{\gamma}{2}t}\left[\cos(\Omega t)-i\frac{\Delta+2U|\Psi_{SS}|^{2}}{\Omega}\sin(\Omega t)\right],$ (53)
$v(t)=e^{-\frac{\gamma}{2}t}\left[-i\frac{U\Psi_{SS}^{2}}{\Omega}\sin(\Omega t)\right],$
$v^{*}(t)=e^{-\frac{\gamma}{2}t}\left[i\frac{(U\Psi_{SS}^{2})^{*}}{\Omega}\sin(\Omega t)\right],$
$u^{*}(t)=e^{-\frac{\gamma}{2}t}\left[\cos(\Omega t)+i\frac{\Delta+2U|\Psi_{SS}|^{2}}{\Omega}\sin(\Omega t)\right].$
While in unstable/parametric regime ($\det(\mathbf{M})<0$) where the parametric gain dominates ($(\Delta+2U|\Psi_{SS}|^{2})^{2}<|U\Psi_{SS}^{2}|^{2}$), we define the growth rate $\Lambda=\sqrt{-\det(\mathbf{M})}=\sqrt{|U\Psi_{SS}^{2}|^{2}-(\Delta+2U|\Psi_{SS}|^{2})^{2}}$. In this case, $\mathbf{M}^{2}=\Lambda^{2}\mathbf{I}$. The Taylor expansion yields a hyperbolic solution:
$e^{\mathbf{M}t}$ $=\sum_{k=0}^{\infty}\frac{t^{2k}\Lambda^{2k}}{(2k)!}\mathbf{I}+\sum_{k=0}^{\infty}\frac{t^{2k+1}\Lambda^{2k}}{(2k+1)!}\mathbf{M}$ (54)
$=\cosh(\Lambda t)\mathbf{I}+\frac{\sinh(\Lambda t)}{\Lambda}\mathbf{M}.$ (55)
This corresponds to the exponential growth of fluctuations (parametric amplification) which can lead to instability if the gain $\Lambda$ exceeds the damping rate $\gamma/2$.
Since $\mathbf{L}_{\text{eff}}$ is time-independent (as a function of constants $\Psi_{SS},\Delta,U,\gamma$), thus the propagator $e^{\mathbf{L}_{\text{eff}}t}$ is defined by the power series of the constant matrix $\mathbf{L}_{\text{eff}}$ scaled by the scalar $t$, and it is a linear time translation invariant system. Next we focus on the noise accumulation integral. The inhomogeneous integral term provides the required drift ($\mathbf{L}_{\text{eff}}$) and the instantaneous noise ($\bm{\xi}$) to satisfy the inhomogeneous Langevin equation (which is the Bogoliubov-de Gennes linear equation describing the fluctuation)
$\frac{d}{dt}\begin{pmatrix}\delta\hat{a}(t)\\ \delta\hat{a}^{\dagger}(t)\end{pmatrix}=\mathbf{L}_{eff}\begin{pmatrix}\delta\hat{a}(t)\\ \delta\hat{a}^{\dagger}(t)\end{pmatrix}+\begin{pmatrix}\xi(t)\\ \xi^{{\dagger}}(t)\end{pmatrix}$ (56)
the solution is
$\begin{pmatrix}\delta\hat{a}(t)\\ \delta\hat{a}^{\dagger}(t)\end{pmatrix}=e^{\mathbf{L}_{eff}t}\begin{pmatrix}\delta\hat{a}(0)\\ \delta\hat{a}^{\dagger}(0)\end{pmatrix}+\int_{0}^{t}e^{\mathbf{L}_{eff}(t-\tau)}\begin{pmatrix}\xi(\tau)\\ \xi^{{\dagger}}(\tau)\end{pmatrix}d\tau,$ (57)
where the first temr is homogeneous solution and the second term is inhomogeneous solution. Note that using Leibniz integral rule, $\frac{d}{dt}\left(\int_{a(t)}^{b(t)}f(t,\tau)d\tau\right)=f(t,b(t))\cdot b^{\prime}(t)-f(t,a(t))\cdot a^{\prime}(t)+\int_{a(t)}^{b(t)}\frac{\partial}{\partial t}f(t,\tau)d\tau$, we have
$\frac{d}{dt}\int_{0}^{t}e^{\mathbf{L}_{eff}(t-\tau)}\begin{pmatrix}\xi(\tau)\\ \xi^{{\dagger}}(\tau)\end{pmatrix}d\tau=\mathbf{L}_{eff}\int_{0}^{t}e^{\mathbf{L}_{eff}(t-\tau)}\begin{pmatrix}\xi(\tau)\\ \xi^{{\dagger}}(\tau)\end{pmatrix}d\tau+\begin{pmatrix}\xi(t)\\ \xi^{{\dagger}}(t)\end{pmatrix}$, (58)
where $\mathbf{L}_{eff}(t-\tau)$ is the result of integral in the exponent $\int_{\tau}^{t}\mathbf{L}_{\text{eff}}dt^{\prime}=\mathbf{L}_{\text{eff}}\cdot(t-\tau)$, in constract to the time-ordered exponential $\mathcal{T}\exp\left(\int_{\tau}^{t}\mathbf{L}_{\text{eff}}(t^{\prime})dt^{\prime}\right)$.
Using the Bogoliubov structure $e^{\mathbf{L}_{eff}t}=\begin{pmatrix}u(t)&v(t)\\ v^{*}(t)&u^{*}(t)\end{pmatrix}$ and $e^{\mathbf{L}_{eff}(t-\tau)}=\begin{pmatrix}u(t-\tau)&v(t-\tau)\\ v^{*}(t-\tau)&u^{*}(t-\tau)\end{pmatrix}$, we have the following noise accumulate integral
$\begin{pmatrix}\hat{\mathcal{F}}_{1}(t)\\ \hat{\mathcal{F}}_{2}(t)\end{pmatrix}=\int_{0}^{t}\begin{pmatrix}u(t-\tau)&v(t-\tau)\\ v^{*}(t-\tau)&u^{*}(t-\tau)\end{pmatrix}\begin{pmatrix}\hat{\xi}(\tau)\\ \hat{\xi}^{\dagger}(\tau)\end{pmatrix}d\tau$ (59)
where $u(t-\tau)\hat{\xi}(\tau)$ and $u^{*}(t-\tau)\hat{\xi}^{{\dagger}}(\tau)$ represent the normal noise that injected at time $\tau$ propagates to time $t$ behaving like a regular particle. $v(t-\tau)\hat{\xi}^{\dagger}(\tau)$ and $v^{*}(t-\tau)\hat{\xi}(\tau)$ represent the anomalous noise. Due to nonlinear pumping, the hole noise (and particle noise) injected at time $\tau$ is converted and mixed into the particle’s (and hole’s) evolution, respectively. This term is essential for the squeezing.
Performing the matrix multiplication explicitly:
$\hat{\mathcal{F}}_{1}(t)$ $=\int_{0}^{t}\left[u(t-\tau)\hat{\xi}(\tau)+v(t-\tau)\hat{\xi}^{\dagger}(\tau)\right]d\tau$ (60)
$\hat{\mathcal{F}}_{2}(t)$ $=\int_{0}^{t}\left[v^{*}(t-\tau)\hat{\xi}(\tau)+u^{*}(t-\tau)\hat{\xi}^{\dagger}(\tau)\right]d\tau$
where $(\hat{\mathcal{F}}_{1}(t))^{\dagger}=\hat{\mathcal{F}}_{2}(t)$. Thus $\hat{\mathcal{F}}_{2}$ represents the fluctuations of the creation operator $\delta\hat{a}^{\dagger}$.
At zero-temperature (vacuum state), we use the commutation relation $\hat{\xi},\hat{\xi}^{\dagger}]=\gamma\delta$ and $\langle\hat{\xi}\hat{\xi}^{\dagger}\rangle=\langle\hat{\xi}^{\dagger}\hat{\xi}\rangle+\langle[\hat{\xi},\hat{\xi}^{\dagger}]\rangle=0+\gamma\delta$, the Markovian noise satisfies the relations
$\langle\hat{\xi}(\tau)\hat{\xi}^{\dagger}(\tau^{\prime})\rangle=\gamma\delta(\tau-\tau^{\prime}),\qquad\langle\hat{\xi}^{\dagger}(\tau)\hat{\xi}(\tau^{\prime})\rangle=\langle\hat{\xi}(\tau)\hat{\xi}(\tau^{\prime})\rangle=\langle\hat{\xi}(\tau)\rangle=0.$ (61)
Thus the particle number fluctuations ($\langle\hat{\mathcal{F}}_{1}^{\dagger}\hat{\mathcal{F}}_{1}\rangle$) corresponds to the normal-ordered correlation $\langle\delta\hat{a}^{\dagger}\delta\hat{a}\rangle$.
$\langle\hat{\mathcal{F}}_{1}^{\dagger}\hat{\mathcal{F}}_{1}\rangle$ $=\int_{0}^{t}d\tau_{1}\int_{0}^{t}d\tau_{2}\langle\hat{\mathcal{F}}_{2}(\tau_{1})\hat{\mathcal{F}}_{1}(\tau_{2})\rangle$ (62)
$=\int_{0}^{t}d\tau_{1}\int_{0}^{t}d\tau_{2}\langle[v^{*}(t-\tau_{1})\hat{\xi}_{1}+u^{*}(t-\tau_{1})\hat{\xi}_{1}^{\dagger}][u(t-\tau_{2})\hat{\xi}_{2}+v(t-\tau_{2})\hat{\xi}_{2}^{\dagger}]\rangle$
$=\int_{0}^{t}d\tau_{1}\int_{0}^{t}d\tau_{2}v^{*}(t-\tau_{1})v(t-\tau_{2})\langle\hat{\xi}(\tau_{1})\hat{\xi}^{\dagger}(\tau_{2})\rangle=\gamma\int_{0}^{t}|v(t-\tau)|^{2}d\tau$
where $\langle\hat{\xi}(\tau_{1})\hat{\xi}^{\dagger}(\tau_{2})\rangle=\gamma\delta(\tau_{1}-\tau_{2})$ the last step is because only the term containing $\langle\hat{\xi}_{1}\hat{\xi}_{2}^{\dagger}\rangle$ survives. Using $|v(t)|=e^{-\gamma t/2}\frac{|U\Psi_{SS}^{2}|}{\Omega}|\sin(\Omega t)|$, we have
$\gamma\int_{0}^{t}|v(t-\tau)|^{2}d\tau$ $=\frac{\gamma|P|^{2}}{\Omega^{2}}\int_{0}^{t}e^{-\gamma(t-\tau)}\sin^{2}(\Omega(t-\tau))d\tau$ (63)
$=\frac{\gamma|P|^{2}}{\Omega^{2}}e^{-\gamma t}\int_{0}^{t}e^{\gamma\tau}\sin^{2}(\Omega(t-\tau))d\tau$
$=\frac{\gamma|P|^{2}}{\Omega^{2}}e^{-\gamma t}\int_{0}^{t}e^{\gamma\tau}\frac{1-\cos(2\Omega(t-\tau))}{2}d\tau$
$=\frac{\gamma|P|^{2}}{\Omega^{2}}e^{-\gamma t}\left[\frac{1}{2}\int_{0}^{t}e^{\gamma\tau}d\tau-\frac{1}{2}\int_{0}^{t}e^{\gamma\tau}\cos(2\Omega(t-\tau))d\tau\right]$
$=\frac{\gamma|P|^{2}}{\Omega^{2}}e^{-\gamma t}\cdot\frac{1}{2}\left[\frac{e^{\gamma t}-1}{\gamma}-\text{Re}\left(\frac{e^{\gamma t}-e^{i2\Omega t}}{\gamma-2i\Omega}\right)\right]$
$=\frac{|P|^{2}}{2\Omega^{2}}\left[(1-e^{-\gamma t})-\gamma\cdot\text{Re}\left(\frac{1-e^{-(\gamma-2i\Omega)t}}{\gamma-2i\Omega}\right)\right]$
The hole fluctuations $\langle\hat{\mathcal{F}}_{2}^{\dagger}\hat{\mathcal{F}}_{2}\rangle$ corresponds to the anti-normal correlation $\langle\delta\hat{a}\delta\hat{a}^{\dagger}\rangle=\langle\hat{\mathcal{F}}_{1}\hat{\mathcal{F}}_{1}^{\dagger}\rangle$,
$\langle\hat{\mathcal{F}}_{1}\hat{\mathcal{F}}_{1}^{\dagger}\rangle$ $=\int_{0}^{t}d\tau_{1}\int_{0}^{t}d\tau_{2}\langle[u(t-\tau_{1})\hat{\xi}_{1}+v(t-\tau_{1})\hat{\xi}_{1}^{\dagger}][v^{*}(t-\tau_{2})\hat{\xi}_{2}+u^{*}(t-\tau_{2})\hat{\xi}_{2}^{\dagger}]\rangle=\gamma\int_{0}^{t}|u(t-\tau)|^{2}d\tau,$ (64)
where the only surviving term is $\langle\hat{\xi}(\tau_{1})\hat{\xi}^{\dagger}(\tau_{2})\rangle$ multiplied by $u(\tau_{1})u^{*}(\tau_{2})$.
Using the Bogoliubov identity $|u|^{2}-|v|^{2}=e^{-\gamma\tau}$, we have
$\langle\hat{\mathcal{F}}_{2}^{\dagger}\hat{\mathcal{F}}_{2}\rangle-\langle\hat{\mathcal{F}}_{1}^{\dagger}\hat{\mathcal{F}}_{1}\rangle=[\hat{\mathcal{F}}_{1},\hat{\mathcal{F}}_{1}^{{\dagger}}]=\gamma\int_{0}^{t}(|u|^{2}-|v|^{2})d\tau=\gamma\int_{0}^{t}e^{-\gamma\tau}d\tau=1-e^{-\gamma t}$ (65)
Thus $\langle\hat{\mathcal{F}}_{2}^{\dagger}\hat{\mathcal{F}}_{2}\rangle$ is the particle number plus the commutation relation term which decays to 1 (restoring the vacuum commutator) as $t\to\infty$.
VII Truncated Wigner Approximation (TWA)
The truncated Wigner approximation (TWA) provides a semi-classical framework to study the many-body dynamics of the driven-dissipative Bose-Hubbard chain by mapping the Lindblad master equation onto a set of coupled stochastic differential equations (SDEs). In TWA the boson operator $\hat{a}_{l}$ is replaced by complex field $\alpha_{l}$ which can be decomposed into its ensemble average and its stochastic fluctuation $\alpha_{l}(t)=\Psi_{l}(t)+\delta\alpha_{l}(t)$ where $\Psi_{l}(t)=\langle\alpha_{l}(t)\rangle$ is the coherent amplitude and $\delta\alpha_{l}(t)$ represents the quantum fluctuations. The total intensity measured in the Wigner representation is
$\langle|\alpha_{l}|^{2}\rangle=|\Psi_{l}|^{2}+\langle|\delta\alpha_{l}|^{2}\rangle$ (66)
where $|\Psi_{l}|^{2}$ is the coherent Power, $\langle|\delta\alpha_{l}|^{2}\rangle=\langle|\alpha|^{2}\rangle-|\langle\alpha\rangle|^{2}$ is the total variance of $\alpha_{l}$. The physical photon number is then $\langle n_{l}\rangle=|\Psi_{l}|^{2}+(\langle|\delta\alpha_{l}|^{2}\rangle-1/2)$. During a deterministic (ordering) process, the mean value $\Psi_{l}$ grows from zero driven by the external field $F$. $\langle|\delta\alpha_{l}|^{2}\rangle=\int|\alpha|^{2}W(\alpha)d^{2}\alpha=1/2$ if the system is in a coherent state and $\langle n_{l}\rangle=|\Psi_{l}|^{2}$, where $W(\alpha)=\frac{2}{\pi}e^{-2|\alpha|^{2}}$. In chaotic regime (bulk site and drain site), the system nonlinearly scrambles the phase of the drive, effectively restoring the $U(1)$ symmetry locally. The nonlinearity $U$ causes trajectories to diverge exponentially[7] and the variance increases ($\langle|\delta\alpha_{l}|^{2}\rangle>1/2$) due to the incoherent thermal excitations. This leads to the loss of phase coherence and the restoration of $U(1)$ symmetry. Kerr nonlinearity and hopping term are invariant under a global phase shift $\alpha_{l}\to\alpha_{l}e^{i\theta}$, which corresponds to photon number conservation. But the drive term $F(\hat{a}_{1}+\hat{a}_{1}^{\dagger})$ at site 1 breaks the U(1) symmetry by fixing a phase.
The bulk site is only coupled to the drive through hopping $J$. In chaotic regime ($U\gg\gamma$), the nonlinear scrambling of phases washes out the drive’s influence. While in a chaotic NESS, the distribution of $\alpha_{2}$ in phase space ($Re[\alpha]$ vs $Im[\alpha]$) is a ring and thus independent of the phase $\theta={\rm arg}\ \alpha$, signifying the restoration of $U(1)$ symmetry despite the presence of a phase-fixed drive at the boundary. Site 3 also in scrambled phase but the high dissipation rate $\gamma_{3}$ keeps the average amplitude $|\alpha_{3}|$ smaller than in the bulk site.
As shown in Fig.4(a), the external drive $F$ fixes the average phase (determined by the driving laser) and prevents the distribution from closing into a ring. The Kerr nonlinearity $U$ makes the rotation frequency dependent on the amplitude ($|\alpha|^{2}$). The external drive $F$ at Site 1 acts as a phase anchor, fixing $\langle\phi\rangle$. In a nonlinear system, the effective resonance frequency (angular velocity in phase space) depends on the photon number: $\Delta_{\text{eff}}=\Delta+U(|\alpha|^{2}-1)=-\dot{\theta}$. In the phase plane, this means points with a larger amplitude ($|\alpha|$) rotate at a higher angular velocity than points with a smaller amplitude.
In Fig.4(a) we set $U=2.5,F=6$. In site 1 the drive $F$ is strong enough to keep the coherent offset from the origin with broken symmetry. Sites 2 and 3, the nonlinearity $U$ in the bulk is strong enough to scramble the phase information, restoring the $U(1)$ symmetry ($\langle\alpha\rangle=0$). The local nonlinearity at site 2 is strong enough to eliminate the phase information inherited from the drive as long as the phase diffusion rate ($\sim U$) is larger than the coherent transfer rate ($\sim J$). We show the final complex amplitudes $\{\alpha_{l}^{(m)}\}$ for a large ensemble of $M=800$ trajectories. We use a Monte-Carlo sampling that reconstructs the steady-state local Wigner function $W_{l}(\alpha)$. The transition from a nonsymmetry distribution in site 1 to a symmetry distribution in site 2 and site 3 confirms the system’s transition from a driven-coherent state to a many-body chaotic state where phase coherence is lost due to the scrambling but energy (particle number) is transported and the system has thermalized and become chaotic locally, restoring the $U(1)$ symmetry $\langle\alpha\rangle=0$. The phase $\theta=\text{arg}(\alpha)$ becomes uniformly distributed over $0,2\pi]$. The vanishing expectation value $\langle\alpha\rangle$ can be the Wigner-weighted integral
$\langle\alpha\rangle=\int\alpha W(\alpha)d^{2}\alpha=\int_{0}^{\infty}\int_{0}^{2\pi}re^{i\theta}W(r,\theta)rdrd\theta=0.$ (67)
where $\alpha=re^{i\theta}$ is the Weyl symbol of $\hat{a}$. In the chaotic regime with ring-shape phase distribution, the distribution $W(r,\theta)$ becomes phase-independent ($W(r)$), and since $\int_{0}^{2\pi}e^{i\theta}d\theta=0$, the integral vanishes. Although the average field $\langle\alpha\rangle$ is zero, the points are distributed far from the origin ($r\gg 0$). The photon number $\langle n\rangle=\langle|\alpha|^{2}\rangle-1/2$ remains large, representing a phase-scrambled high-density state. In (b) we set $U=0,F=1$. In this linear limit where the coherent information is preserved, the system behaves as a collection of driven-damped harmonic oscillators. The ring-shape distribution observed are the circular uncertainty clouds of vacuum noise with $1/2$ width centered at the coherent amplitude $\alpha_{SS}$. The trajectories form a Gaussian cloud centered at the origin with a width of $1/2$. This represents the vacuum state. The drift away from the origin signifies broken $U(1)$ symmetry due to the drive. In TWA, the vacuum state has variance $\langle|\alpha|^{2}\rangle=\frac{1}{2}$ as a result of the commutation relation $\hat{a},\hat{a}^{\dagger}]=1$, play the role of quantum noise and zero-point fluctuations which allows the system to explore the phase space and trigger non-linear effects or chaos[2].
The Monte Carlo sampling in TWA applies $\langle\alpha\rangle=\frac{1}{M}\sum_{m=1}^{M}\alpha^{(m)}$. As $M$ increases, the random phases of the trajectories $\alpha^{(m)}$ cancel out pairwise in the complex plane, leading to $\langle\alpha\rangle\to 0$. The TWA works by mapping the quantum master equation to a set of mutually independent stochastic differential equations (Langevin equations). The sampling start with $M$ copies of the system by using an ensemble of $M$ classical-like trajectories. Each copy’s initial state $\alpha(0)$ is sampled from the Wigner distribution of the vacuum (a Gaussian cloud of width $1/\sqrt{2}$). At $t=0$, the vacuum state $|0\rangle$ in Wigner representation is a Gaussian with variance $1/2$,
$\alpha_{l}^{(m)}(0)=\frac{1}{\sqrt{2}}(\eta_{1,l}^{(m)}+i\eta_{2,l}^{(m)}),\quad\eta\sim\mathcal{N}(0,1).$ (68)
Each trajectory is evolved independently using the Heun integrator. The stochastic noise term $\sqrt{\gamma/2}dW$ simulates the quantum fluctuations entering from the environment. Each trajectory $m$ evolves according to a stochastic differential equation (Langevin equation) $\alpha(t+dt)=\alpha(t)+\mathcal{A}(\alpha)dt+\sqrt{\frac{\gamma}{2}}d\mathcal{W}$. The noise $d\mathcal{W}$ ensures that the diffusion of trajectories satisfy the uncertainty principle. Because the underlying Fokker-Planck equation (FPE) describes the evolution of the probability density $W(\alpha,t)$, the set of $M$ points at any time $t$ serves as a statistical histogram of the quantum state in phase space. As $M\to\infty$, the density of these points perfectly reconstructs the continuous density $W(\alpha,t)$ defined by the FPE.
The symmetric (Weyl) ordering[16] of the Hamiltonian leads to the following correspondence for the Kerr nonlinearity
$\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}=\hat{n}(\hat{n}-1)=\hat{n}^{2}-\hat{n}\to(|\alpha|^{2}-\frac{1}{2})^{2}-\frac{1}{4}-(|\alpha|^{2}-\frac{1}{2})$ (69)
$=(|\alpha|^{4}-|\alpha|^{2})-(|\alpha|^{2}-1/2)=|\alpha|^{4}-2|\alpha|^{2}+\frac{1}{2}.$
This results in a renormalized nonlinearity in the drift equations. The quantum fluctuations are incorporated through the initial conditions and the stochastic noise terms derived from the dissipators. In the Wigner-Weyl correspondence, the classical variable $|\alpha|^{2}$ does not map to particle number $\hat{n}=\hat{a}^{\dagger}\hat{a}$ but maps to the symmetrically ordered product $|\alpha|^{2}=\frac{1}{2}(\hat{a}^{\dagger}\hat{a}+\hat{a}\hat{a}^{\dagger})=\hat{a}^{\dagger}\hat{a}+\frac{1}{2}[\hat{a},\hat{a}^{\dagger}]=\hat{n}+\frac{1}{2}$. Thus the averaged population can be obtained by the stochastic average of $|\alpha|^{2}$ subtract a $1/2$ quantum bias, $\langle n\rangle=\overline{|\alpha|^{2}}-1/2$, and the particle number operator in Wigner space is projected as $\hat{a}^{\dagger}\hat{a}\to|\alpha|^{2}-\frac{1}{2}$. At vacuum state ($|0\rangle$) where $\langle\hat{n}\rangle=0$, the variance of the complex field is exactly 1/2, i.e., the minimal uncertainty state. Here the minimal variance $1/2$ for the complex variable $\alpha=X+iP$ directly corresponds to the Heisenberg minimal uncertainty state. For commutator $\hat{X},\hat{P}]=\frac{i}{2}$ with $\hat{X}=\frac{\hat{a}+\hat{a}^{\dagger}}{2}={\rm Re}\alpha$ and $\hat{P}=\frac{\hat{a}-\hat{a}^{\dagger}}{2i}={\rm Im}\alpha$, the Heisenberg uncertainty principle states
$\Delta X\Delta P\geq\frac{1}{2}|\langle[\hat{X},\hat{P}]\rangle|=\frac{1}{4}.$ (70)
For a symmetric vacuum state, the variance of the complex variable is $\langle|\alpha|^{2}\rangle=\langle X^{2}\rangle+\langle P^{2}\rangle=1/2$, with $\langle X^{2}\rangle=\langle P^{2}\rangle=1/4$ and $\Delta X=\Delta P=1/2$ (which is possibel only for the vacuum or a coherent/unsqueezed state). The product of uncertainties in linear system with Gaussian state is $\Delta X\Delta P=\frac{1}{4}$, corresponds to the minimal uncertainty state and minimal variance ($1/2$) in TWA. This is consistent with above symmetric ordering requirement $\overline{|\alpha|^{2}}=\langle\hat{n}\rangle+1/2$. The noise $\sqrt{\gamma/2}dW$ which inject randomness is isotropic while the nonlinearity ($U$) is anisotropic and cause non-Gaussian state with $\Delta X\Delta P>\frac{1}{4}$. As the system relaxes from the vacuum to a driven NESS, the total variance $|\alpha|^{2}$ increases due to the fluctuation part of TWA which suppressing the coherent amplitude and restoring the $U(1)$ symmetry. While the sum of variances is bounded by the noise term, the individual quadrature variance (e.g., $\Delta X^{2}$) can be compressed below 1/4.
The Kerr term $\hat{n}(\hat{n}-1)$ maps to the Wigner symbol $H_{int}=\frac{U}{2}(|\alpha|^{4}-2|\alpha|^{2}+1/2)$. The drift $\mathcal{A}$ in Langevin equation is derived from the derivative (Poisson bracket) $i\frac{\partial H_{int}}{\partial\alpha^{*}}$,
$\frac{\partial}{\partial\alpha^{*}}\left[\frac{U}{2}(|\alpha|^{2}\alpha\alpha^{*}-2\alpha\alpha^{*})\right]=\frac{U}{2}(2|\alpha|^{2}\alpha-2\alpha)=U(|\alpha|^{2}-1)\alpha.$ (71)
where the quantum bias 1-1 is related to the above bias 1/2 in the operator $\alpha$ in symmetric (Weyl) ordering and ensures the classical trajectories obeys the quantum commutation relations during evolution.
The evolution of complex amplitudes $\alpha_{l}$ for a chain of length $L$ is governed by the Langevin equation (in the Itô sense)
$d\alpha_{l}=\dot{\alpha}_{l}=\mathcal{A}_{l}(\{\alpha\})dt+\sum_{k}\mathcal{B}_{lk}d\mathcal{W}_{k}$ (72)
where the first term in right-hand-side is the deterministic drift term and second term is the stochastic diffusion term. The deterministic drift term $\mathcal{A}_{l}$ reads
$\mathcal{A}_{l}=-i\left[\Delta\alpha_{l}+U(|\alpha_{l}|^{2}-1)\alpha_{l}-J(\alpha_{l-1}+\alpha_{l+1})+F\delta_{l,1}\right]-\frac{\gamma_{l}}{2}\alpha_{l}$ (73)
Note that in any cases, we consider the deturning term $\propto\Delta n<0$ (i.e., $i\Delta\alpha$ in Heisenberg equation of motion), in consistent with the content in above sections. From the Heisenberg equation $\dot{\hat{a}}=-i[\hat{a},\hat{H}$], the interaction part is $i[\hat{a},\frac{U}{2}\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}]=-iU\hat{a}^{\dagger}\hat{a}\hat{a}=-iU\hat{n}\hat{a}$. The TWA variables $\alpha$ in symmetric (Weyl) ordering can be used to map the operator product $\hat{n}\hat{a}=\hat{a}^{\dagger}\hat{a}\hat{a}$ to phase space, by the Weyl symbol of $\hat{a}^{\dagger}\hat{a}\hat{a}$ $|\alpha|^{2}\alpha-\alpha=(|\alpha|^{2}-1)\alpha$, where 1-1 is a quantum correction (ordering bias) that accounts for the commutation relation $\hat{a},\hat{a}^{\dagger}]=1$. The term $|\alpha|^{2}\alpha$ represents the classical mean-field interaction. The term $\alpha$ represents the self-interaction with vacuum noise that subtracts the over-counted zero-point energy inherent in the Wigner representation. The hopping term $J(\alpha_{l-1}+\alpha_{l+1})$ describes the coherent exchange of particles between neighboring sites in the chain. The dissipative damping $\frac{\gamma_{l}}{2}\alpha_{l}$ comes from the $\mathcal{D}[\sqrt{\gamma}\hat{a}$] term in the Lindblad equation. It causes the amplitude to decay exponentially toward zero in the absence of a drive. The stochastic term represents the vacuum noise entering through the loss channels, with the Markovian white noise $\langle d\mathcal{W}_{l}(t)d\mathcal{W}_{k}^{*}(t^{\prime})\rangle=\delta_{lk}\delta(t-t^{\prime})dt$. Since the loss is local to each site, $\mathcal{B}$ is a diagonal matrix with $\mathcal{B}_{ll}=\sqrt{\gamma_{l}/2}$ and scales the random noise by the square root of the dissipation rate. The complex wiener process $d\mathcal{W}_{l}$ represents independent random variables sampled at each time step. The evolution of the fluctuation spinor $\mathbf{v}=(\delta\hat{a},\delta\hat{a}^{\dagger})^{T}$ is governed by the linearized equation $\dot{\mathbf{v}}=\mathbf{L}_{\text{eff}}\mathbf{v}$. The matrix $\mathbf{L}_{\text{eff}}$ is defined as the Jacobian matrix of the total deterministic drift function $\dot{\alpha}=\mathcal{A}(\alpha,\alpha^{*})$,
$\mathbf{L}_{\text{eff}}=\begin{pmatrix}\frac{\partial\dot{\alpha}}{\partial\alpha}&\frac{\partial\dot{\alpha}}{\partial\alpha^{*}}\\ \frac{\partial\dot{\alpha}^{*}}{\partial\alpha}&\frac{\partial\dot{\alpha}^{*}}{\partial\alpha^{*}}\end{pmatrix}$ (74)
Based on the stochastic differential equation
$d\alpha_{l}=\left[-i(\Delta\alpha_{l}+U(|\alpha_{l}|^{2}-1)\alpha_{l}-J\sum_{k=l\pm 1}\alpha_{k}+F\delta_{l,1})-\frac{\gamma_{l}}{2}\alpha_{l}\right]dt+\sqrt{\frac{\gamma_{l}}{2}}d\mathcal{W}_{l}(t),$ (75)
the deterministic drift term for a site is $\dot{\alpha}=-i\Delta\alpha-iU(\alpha^{2}\alpha^{*}-\alpha)+iJ\sum\alpha_{k}-iF-\frac{\gamma}{2}\alpha$.
$L_{11}=\frac{\partial\dot{\alpha}}{\partial\alpha}=-i\Delta-iU(2|\alpha|^{2}-1)-\frac{\gamma}{2}=-i(\Delta+2U|\alpha|^{2}-U)-\frac{\gamma}{2},$ (76)
$L_{12}=\frac{\partial\dot{\alpha}}{\partial\alpha^{*}}=-iU\alpha^{2},$
$L_{21}=\frac{\partial\dot{\alpha}^{*}}{\partial\alpha}=iU(\alpha^{*})^{2},$
$L_{22}=\frac{\partial\dot{\alpha}^{*}}{\partial\alpha^{*}}=i\Delta+iU(2|\alpha|^{2}-1)-\frac{\gamma}{2}=i(\Delta+2U|\alpha|^{2}-U)-\frac{\gamma}{2}.$
Thus the $\mathbf{L}_{\text{eff}}$ in Eq.(48) can be rewritten as
$\mathbf{L}_{\text{eff}}=\begin{pmatrix}-i(\Delta+2U|\alpha|^{2}-U)-\frac{\gamma}{2}&-iU\alpha^{2}\\ iU(\alpha^{*})^{2}&i(\Delta+2U|\alpha|^{2}-U)-\frac{\gamma}{2}\end{pmatrix}$ (77)
A discussion on quantum metric framework is presented in Appendix.
In linear limit $U=0$, $L_{11}=-i\Delta-\frac{\gamma}{2}=L^{*}_{22}$, The total equation of motion for $\alpha$ (which includes the hopping and drive terms that are treated as constants in the local Jacobian derivative) reads $\dot{\alpha}=(-i\Delta-\frac{\gamma}{2})\alpha+iJ\sum\alpha_{k}-iF$. For a single driven resonator ($J=0$), $\alpha=\frac{iF}{-i\Delta-\frac{\gamma}{2}}$ and $|\alpha|=\frac{|F|}{\sqrt{\Delta^{2}+\frac{\gamma^{2}}{4}}}$. The center of phase space ($\text{Re}[\alpha]=\text{Im}[\alpha]=0$) represents the vacuum state with zero photon and zero drive. A finite coherent drive $F$ breaks U(1) symmetry and causes a displacement $\alpha_{SS}$ where $|\alpha_{SS}|$ shows coherent photons number and $\text{arg}(\alpha_{SS})$ shows the phase of the light that is locked to the phase of the driving laser. The steady-state displacement $\alpha_{l}$ is found by solving $\dot{\alpha}_{l}=0$. Define the complex impedance $z_{l}=-\Lambda_{l}$, we have
$z_{1}\alpha_{1}-iJ\alpha_{2}$ $=-iF$ (78)
$z_{2}\alpha_{2}-iJ(\alpha_{1}+\alpha_{3})$ $=0$
$z_{3}\alpha_{3}-iJ\alpha_{2}$ $=0$
Similar to Eq.(26), we have
$\alpha_{1}=\frac{-iF}{z_{1}+\frac{J^{2}}{z_{2}+J^{2}/z_{3}}},$ (79)
$\alpha_{2}=\left(\frac{iJ}{z_{2}+J^{2}/z_{3}}\right)\alpha_{1},$
$\alpha_{3}=\left(\frac{iJ}{z_{3}}\right)\alpha_{2}$
where $J\alpha_{1}$ and $J\alpha_{2}$ play the role of effective drive that allows the coherence propagation through the chain. In phase space, this means the centers of the Wigner distributions for all three sites are shifted away from $(0,0)$.
$|\alpha_{3}|=\frac{|FJ^{2}|}{|z_{1}z_{2}z_{3}+J^{2}(z_{1}+z_{3})|},$ (80)
$\frac{|\alpha_{3}|}{|\alpha_{2}|}=\frac{|J|}{\sqrt{\Delta^{2}+\frac{\gamma_{3}^{2}}{4}}},$
$\frac{|\alpha_{2}|}{|\alpha_{1}|}=\frac{|J|\sqrt{\Delta^{2}+\frac{\gamma_{3}^{2}}{4}}}{\sqrt{\left(J^{2}-\Delta^{2}+\frac{\gamma_{2}\gamma_{3}}{4}\right)^{2}+\Delta^{2}\frac{(\gamma_{2}+\gamma_{3})^{2}}{4}}}$
Despite the absence of $F$ in site 2 and site 3, all sites are connected by the hopping $J$: Since $\alpha_{1}\propto F$, and since $\alpha_{2}$ is proportional to $\alpha_{1}$, and $\alpha_{3}$ is proportional to $\alpha_{2}$, all site amplitudes are linearly scaled by $F$, and the energy and the specific phase of the drive will transport through $J$. This is why in Fig.4(b), all three sites are drifted away from the origin. They have all inherited a portion of the drive’s coherent phase and amplitude. We set $\Delta=-1$ to consider the case that the pump frequency is higher than the cavity frequency. As long as $|\Delta|<1$, the bulk site has larger steady state amplitude than the drive site.
Figure 4: (a) The chaotic regime ($U=2.5,F=6,J=0.8$) where nonlinearity introduces phase scrambling. At site 1, the drive $F$ is strong enough to maintain a coherent offset. At sites 2 and 3, the scrambling $U$ is stronger than the transfer $J$. This generates diffusion and randomizes the phase, and the Wigner distribution is centered at (0,0), effectively restoring the $U(1)$ symmetry. (b) The linear regime ($U=0,F=1,J=0.8$) where all sites are displaced from the origin. The phase is pinned and U(1) symmetry is broken by the drive because the coherence $J$ is dominate over the randomizing force.
[…]
Figure 6: Amplitude at each site as a function of coupling ($J$) and the external driving strength ($F$), with the detuning fixed at $\Delta=-1$. In all three sites, the amplitude increases linearly as the driving force $F$ increases. There is a distinct vertical bright region in all sites, most notably around $J\approx 0.75-1.0$. This represents the optimal coupling strength where the external driving frequency matches the system’s normal modes. $|\alpha_{1}|$ and $|\alpha_{2}|$ reach similar maximum values (around $3.5$ at $F=2.0$). $|\alpha_{3}|$ has a significantly lower maximum amplitude (around $1.75$ at $F=2.0$). This is consistent with site 3 having a much higher decay rate ($\gamma_{3}=2.0$), which drains the energy faster than it can accumulate.
In the absence of quantum noise (classical mean-field limit or deterministic limit), the stochastic differential equation is reduced to the deterministic ordinary differential equation
$\frac{d\alpha_{l}}{dt}=\mathcal{A}_{l}(\{\alpha\})=-i\left[\Delta\alpha_{l}+U(|\alpha_{l}|^{2}-1)\alpha_{l}-J(\alpha_{l-1}+\alpha_{l+1})+F\delta_{l,1}\right]-\frac{\gamma_{l}}{2}\alpha_{l}$ (81)
This is the damped Gross-Pitaevskii Equation. The presence of the 1-1 in the interaction term is a "quantum remnant" of the symmetric ordering used in Wigner-Weyl calculus, distinguishing it from purely classical physics. Note that the drift term $\mathcal{A}_{l}$ in TWA contains quantum bias inside the nonlinearity $U(|\alpha_{l}|^{2}-1)\alpha_{l}$. While a purely classical GPE would typically use $U|\alpha_{l}|^{2}\alpha_{l}$. The TWA drift accounts for the fact that even in the average motion, the particles are interacting with their own symmetric vacuum energy. The initial fluctuations can be sampled from $\alpha_{l}(0)\sim\mathcal{N}(0,1/2)+i\mathcal{N}(0,1/2)$, which ensures the uncertainty required by Heisenberg principle even at $t=0$. While the stochastic term $\sum_{k}\mathcal{B}_{lk}d\mathcal{W}_{k}$ cause the quantum noise entering from the environment and leads to dynamical fluctuations, which prevents the dissipative system to lose quantum uncertainty and collapse to a purely classical state with the deterministic drift $\mathcal{A}_{l}$. The stochastic term also ensuring that the field $\alpha$ maintains a minimum variance of $1/2$ (the vacuum noise) even as it relaxes toward NESS. Whether that NESS is coherent or chaotic depends on whether the variance stays at $1/2$ or blows up due to the nonlinear scrambling of those trajectories.
A challenge in TWA for nonlinear systems is the divergence of trajectories for large $|\alpha_{l}|$. The cubic term $U|\alpha_{l}|^{2}\alpha_{l}$ can lead to numerical overflows. We implement a fixed-step Heun’s method (a predictor-corrector algorithm for SDEs) to enhance stability and utilize a sufficiently small time step $\Delta t$ to ensure that the approximation captures the competition between parametric gain and dissipation accurately.
The Wigner-Weyl correspondence leads to the following drift equation for site $l$:
$i\frac{\partial\alpha_{l}}{\partial t}=-f(\alpha_{l})-J\sum_{k\in\text{neigh}}\alpha_{k}+F\delta_{l,1}-i\frac{\gamma_{l}}{2}\alpha_{l}$
where $f(\alpha_{l})=-\Delta\alpha_{l}-U(|\alpha_{l}|^{2}-1)\alpha_{l}$. Note the 1-1 correction in the Kerr term, which arises from the symmetric ordering of the operators. The single-photon loss at rate $\gamma_{l}$ introduces Gaussian white noise $\xi_{l}(t)$. The full SDE (Langevin equation) in the Itô sense is
$d\alpha_{l}=\left[-i(\Delta\alpha_{l}+U(|\alpha_{l}|^{2}-1)\alpha_{l}-J\sum_{k}\alpha_{k}+F\delta_{l,1})-\frac{\gamma_{l}}{2}\alpha_{l}\right]dt+\sqrt{\frac{\gamma_{l}}{2}}d\mathcal{W}_{l}(t)$ (82)
where $d\mathcal{W}_{l}$ is a complex Wiener process such that $\langle d\mathcal{W}_{l}d\mathcal{W}_{k}^{*}\rangle=\delta_{lk}dt$. The diffusion coefficient $\sqrt{\gamma/2}$ appears in the stochastic update step of the Langevin equation is the pre-factor of the random noise (Wiener process) $dW$, $\alpha(t+\Delta t)=\alpha_{drift}+\sqrt{\frac{\gamma}{2}}d\mathcal{W}$, reflecting the effect from environment that balance the damping term ($\gamma/2$) and maintaining the uncertainty principle.
Algorithm 2 Stabilized TWA for Driven-Dissipative NESS
1:  Input: Parameters $U,J,\Delta,\gamma,F$, time step $\Delta t$, total time $T_{max}$, ensemble size $M$.
2:  Initialize:
3:     Create an ensemble of $M$ trajectories.
4:
5:  for each trajectory $m=1$ to $M$ do
6:         Sample $\alpha_{l}^{(m)}(0)\sim\mathcal{N}(0,1/2)+i\mathcal{N}(0,1/2)$ for all sites $l\in\{1,\dots,L\}$.
7:
8:  end for
9:  Time Evolution (Loop $t=0$ to $T_{max}$ in steps of $\Delta t$):
10:  for each trajectory $m=1$ to $M$ do
11:   1. Generate Noise:
12:      Sample complex Gaussian noise $dW_{l}$ such that $\text{Var}(\text{Re},\text{Im})=\Delta t/2$.
13:   2. Predictor Step (Euler):
14:      $\text{hop}_{l}=\alpha_{l-1}+\alpha_{l+1}$
15:      $\mathcal{A}_{l}(\alpha)=-i\left[\Delta\alpha_{l}+U(|\alpha_{l}|^{2}-1)\alpha_{l}-J\cdot\text{hop}_{l}+F\delta_{l,1}\right]-\frac{\gamma_{l}}{2}\alpha_{l}$
16:      $\alpha_{l}^{temp}=\alpha_{l}(t)+\mathcal{A}_{l}(\alpha(t))\Delta t+\sqrt{\gamma_{l}/2}dW_{l}$
17:   3. Corrector Step (Heun/RK2):
18:      $\alpha_{l}(t+\Delta t)=\alpha_{l}(t)+\frac{1}{2}\left[\mathcal{A}_{l}(\alpha(t))+\mathcal{A}_{l}(\alpha^{temp})\right]\Delta t+\sqrt{\gamma_{l}/2}dW_{l}$
19:   4. Stability Check:
20:      IF $\max|\alpha|>\text{Threshold}$ THEN discard trajectory.
21:  end for
22:  Observables (Ensemble Average):
23:     Compute mean occupation: $\langle n_{l}(t)\rangle=\frac{1}{M_{valid}}\sum_{m}(|\alpha_{l}^{(m)}(t)|^{2}-1/2)$.
24:     Compute coherence: $\langle a_{l}^{\dagger}a_{k}\rangle=\frac{1}{M}\sum\alpha_{l}^{*}\alpha_{k}-\frac{1}{2}\delta_{lk}$.
25:  Output: Time-dependent population and steady-state correlation functions.
Thus the quantum bias 1-1 term in the drift and the 1/2-1/2 term in the occupation calculation are crucial to account for the commutation relations in the Wigner representation. The Heun’s metho using a predictor-corrector scheme for stochastic convergence in the presence of $U|\alpha|^{2}\alpha$ nonlinearities. The vacuum Noise term $\sqrt{\gamma/2}$ ensures that the system satisfies the fluctuation-dissipation theorem, allowing it to relax to the correct NESS.
VIII Quantum metric framework
For this time-independent linearized stability non-Hermitian matrix $\mathbf{L}_{\text{eff}}$ in Eq.(77), we can decomposition it into Pauli basis
$i\mathbf{L}_{\text{eff}}=\frac{-i\gamma}{2}\mathbf{I}+\frac{1}{2}\sum_{j=x,y,z}(\kappa_{j}+i\lambda_{j})\sigma_{j},$ (83)
where $\vec{\kappa}=(0,-2U\text{Im}(\alpha^{2}),2\Omega)$ and $\vec{\lambda}=(0,2U\text{Re}(\alpha^{2}),0)$. The metric operator $\rho(t)$[9, 10, 15], which is Hermitian positive-definite and time-dependent to restore unitarity and preserve the normalization of fluctuations, has the form $\rho(t)=A(t)\mathbf{I}+(\zeta_{1}(t)\vec{\kappa}+\zeta_{2}(t)\vec{\lambda}+\zeta_{3}(t)(\vec{\kappa}\times\vec{\lambda}))\cdot\vec{\sigma}$ with $A(t)=\frac{1}{2}\text{Tr}(\rho(t))$. Since $i\mathbf{L}_{\text{eff}}^{{\dagger}}\rho(t)-\rho(t)i\mathbf{L}_{\text{eff}}=i\dot{\rho}(t)$, there is a set of coupled differential equations
$\frac{d\zeta_{1}}{dt}=\zeta_{3}(\vec{\kappa}\cdot\vec{\lambda})=0,$ (84)
$\frac{d\zeta_{2}}{dt}=-A(t)-|\vec{\kappa}|^{2}\zeta_{3}(t),$
$\frac{d\zeta_{3}}{dt}=\zeta_{2}(t),$
$\frac{dA}{dt}=-\zeta_{1}(\vec{\kappa}\cdot\vec{\lambda})-|\vec{\lambda}|^{2}\zeta_{2}(t)=-|\vec{\lambda}|^{2}\zeta_{2}(t),$
which can be solved as
$\zeta_{1}(t)=\zeta_{1}(0),$ (85)
$\frac{dA}{dt}=-|\vec{\lambda}|^{2}\frac{d\zeta_{3}}{dt},$
$A(t)=A(0)-|\vec{\lambda}|^{2}(\zeta_{3}(t)-\zeta_{3}(0)),$
$\ddot{\zeta}_{3}+(|\vec{\kappa}|^{2}-|\vec{\lambda}|^{2})\zeta_{3}=-A(0)-|\vec{\lambda}|^{2}\zeta_{3}(0).$
The eigenvalues of $\rho(t)$ are positive, thus $A(t)\pm|(\zeta_{1}(t)\vec{\kappa}+\zeta_{2}(t)\vec{\lambda}+\zeta_{3}(t)(\vec{\kappa}\times\vec{\lambda}))|>0$. The eigenvalues of $i\mathbf{L}_{\text{eff}}$ are $\lambda_{\pm}=-i\frac{\gamma}{2}\pm\sqrt{(\Delta+2U|\alpha|^{2}-U)^{2}-U^{2}|\alpha|^{4}}$, with the corresponding eigenvectors $\langle u_{\pm}|=(U\alpha^{2},\pm\sqrt{(\Delta+2U|\alpha|^{2}-U)^{2}-U^{2}|\alpha|^{4}}-(\Delta+2U|\alpha|^{2}-U))$. At exceptional point where $(\Delta+2U|\alpha|^{2}-U)^{2}=U^{2}|\alpha|^{4}$, both the eigenvectors and eigenvalues coalesce at $\lambda=-i\gamma/2$, with $\langle u_{\pm}|=(U\alpha^{2},-(\Delta+2U|\alpha|^{2}-U))$ and $\langle u_{+}|u_{-}\rangle=U^{2}|\alpha|^{4}+(\Delta+2U|\alpha|^{2}-U)^{2}=2U^{2}|\alpha|^{4}$ can be normalized by $\sqrt{\langle u_{+}|u_{+}\rangle}\sqrt{\langle u_{-}|u_{-}\rangle}$, where
$\langle u_{+}|u_{+}\rangle=2(\Delta+2U|\alpha|^{2}-U)^{2}-2(\Delta+2U|\alpha|^{2}-U)\sqrt{(\Delta+2U|\alpha|^{2}-U)^{2}-U^{2}|\alpha|^{4}},$ (86)
$\langle u_{-}|u_{-}\rangle=2(\Delta+2U|\alpha|^{2}-U)^{2}+2(\Delta+2U|\alpha|^{2}-U)\sqrt{(\Delta+2U|\alpha|^{2}-U)^{2}-U^{2}|\alpha|^{4}}.$
For $s=4(\Delta+2U|\alpha|^{2}-U)^{2}-4U^{2}|\alpha|^{4}>0$, in which case the eigenvalues relative to the global decay $i\gamma/2$ are real and thus corresponds to pure oscillations $\zeta_{3}(t)=C_{1}\cos(\sqrt{s}t)+C_{2}\sin(\sqrt{s}t)+C_{3}$, we define $H_{PT}:=i\mathbf{L}_{\text{eff}}+\frac{i\gamma}{2}\mathbf{I}$ and $\rho(t)=e^{\gamma t}\rho_{0}$ with $\rho_{0}$ the geometric core, then the quasi-Hermiticity condition $i\dot{\rho}_{0}=H_{PT}^{{\dagger}}\rho_{0}-\rho_{0}H_{PT}=H_{PT}^{{\dagger}}\rho(t)-\rho(t)H_{PT}=0$ ($(i\mathbf{L}_{\text{eff}})^{\dagger}\rho(t)-\rho(t)(i\mathbf{L}_{\text{eff}})=i\gamma\rho(t)$) signifying a time-independent metric. A static metric implies a static energy landscape[11]. For a positive-definite metric $\rho=(\rho^{1/2})^{{\dagger}}\rho^{1/2}$ (with unique Hermitian square root $\rho^{1/2}$) that meets the quasi-Hermiticity condition $H_{PT}^{{\dagger}}=\rho_{0}H_{PT}\rho_{0}^{-1}=\rho(t)H_{PT}\rho^{-1}(t)$, $H_{PT}$ is similar to a Hermitian operator $\rho^{1/2}_{0}H_{PT}\rho^{-1/2}_{0}$. The eigenvectors in mapped space are $|w_{\pm}\rangle=\rho_{0}^{1/2}|u_{\pm}\rangle$. The orthogonality $\langle w_{+}|w_{-}\rangle=\langle u_{+}|\rho_{0}|u_{-}\rangle=\langle u_{+}|\rho(t)|u_{-}\rangle=0$ where $|w_{\pm}\rangle$ are the eigenvectors of the Hermitian Hamiltonian $\rho_{0}^{1/2}H_{PT}\rho_{0}^{-1/2}$, and Hermitian operators always have orthogonal eigenvectors for distinct eigenvalues. In the $\mathcal{PT}$-symmetric phase with stationary metric, it is constructed to satisfy the biorthogonality condition $\langle u_{m}|\rho|u_{n}\rangle=\delta_{mn}$. The stationary metric can be constructed into Hermitian form $\rho_{0}=\sum_{n=\pm}|l_{n}\rangle\langle l_{n}|$, where $\langle l_{n}|$ are the left eigenvectors of $H_{PT}$ or $i\mathbf{L}_{\text{eff}}$ (also, right eigenvectors of $H_{PT}^{{\dagger}}$ or $i\mathbf{L}_{\text{eff}}$) satisfying $\langle l_{m}|u_{n}\rangle=\delta_{mn}$ (this orthogonality condition is valid only in the PT-symmetric phase where the Hamiltonian is quasi-Hermitian). This construction ensures the metric is the inverse of overlap matrix of the right eigenvectors. While in the PT-broken phase, $\rho=\sum_{n}|l_{n}\rangle\langle l_{n}|$ results in a metric that is not positive-definite and corresponds to a non-unitary mapping. To preserve a consistent probability interpretation in the broken phase, the metric must be time-dependent. The above similar transformation can be reproduced by the Dyson map
$\rho^{1/2}(t)i\mathbf{L}_{\text{eff}}\rho^{-1/2}(t)+i\dot{\rho^{1/2}}(t)\rho^{-1/2}(t)=\rho_{0}^{1/2}H_{PT}\rho_{0}^{-1/2}$ (87)
where $\rho^{1/2}(t)i\mathbf{L}_{\text{eff}}\rho^{-1/2}(t)=(e^{\gamma t/2}\rho_{0}^{1/2})(H_{PT}-i\frac{\gamma}{2}\mathbf{I})(e^{-\gamma t/2}\rho_{0}^{-1/2})=\rho_{0}^{1/2}H_{PT}\rho_{0}^{-1/2}-i\frac{\gamma}{2}\mathbf{I}$, $i\dot{\rho^{1/2}}(t)\rho^{-1/2}(t)=i(\frac{\gamma}{2}\rho^{1/2}(t))\rho^{-1/2}(t)=i\frac{\gamma}{2}\mathbf{I}$.
In $\mathcal{PT}$-Symmetric phase, the time-dependence of metric is only due to the global decay, $i\dot{\rho}(t)=i\gamma\rho(t)$, and the state vector is $|\mathbf{v}(t)\rangle=|\mathbf{v}(t)\rangle=(\delta\hat{a}(t),\delta\hat{a}^{\dagger}(t))^{T}=e^{(-\gamma/2)t}(c_{+}e^{i(-\sqrt{s}/2)t}|u_{+}\rangle+c_{-}e^{i(\sqrt{s}/2)t}|u_{-}\rangle)$, such that
$\delta\hat{a}(t)=e^{-\frac{\gamma}{2}t}(c_{+}e^{-i(\sqrt{s}/2)t}+c_{-}e^{i(\sqrt{s}/2)t})U\alpha^{2},$ (88)
$\delta\hat{a}^{\dagger}(t)=e^{-\frac{\gamma}{2}t}\left[c_{+}e^{-i(\sqrt{s}/2)t}(\sqrt{s}/2-(\Delta+2U|\alpha|^{2}-U))-c_{-}e^{i(\sqrt{s}/2)t}(\sqrt{s}/2+(\Delta+2U|\alpha|^{2}-U))\right],$
$|\delta a|^{2}+|\delta a^{\dagger}|^{2}=e^{-\gamma t}\{|c_{+}|^{2}\left[|U\alpha^{2}|^{2}+(\sqrt{s}/2-(\Delta+2U|\alpha|^{2}-U))^{2}\right]+|c_{-}|^{2}\left[|U\alpha^{2}|^{2}+(\sqrt{s}/2+(\Delta+2U|\alpha|^{2}-U))^{2}\right]$
$+2\text{Re}\left(c_{+}c_{-}^{*}e^{-2i(\sqrt{s}/2)t}\left[|U\alpha^{2}|^{2}-((\sqrt{s}/2)^{2}-(\Delta+2U|\alpha|^{2}-U)^{2})\right]\right)\}$
The Euclidean norm reads
$\langle\mathbf{v}(t)|\mathbf{v}(t)\rangle=e^{-\gamma t}[|c_{+}|^{2}\langle u_{+}|u_{+}\rangle+|c_{-}|^{2}\langle u_{-}|u_{-}\rangle+2\text{Re}(c_{+}^{*}c_{-}e^{2i(\sqrt{s}/2)t}\langle u_{+}|u_{-}\rangle)].$ (89)
Because the eigenvectors are non-orthogonal ($\langle u_{+}|u_{-}\rangle\neq 0$), the Euclidean norm oscillates in time even though the system is stable. The metric framework is used to define a metric that satisfies $\langle\mathbf{v}(t)|\rho_{0}|\mathbf{v}(t)\rangle=e^{-\gamma t}(|c_{+}|^{2}+|c_{-}|^{2})=e^{-\gamma t}$ and $\langle\mathbf{v}(t)|\rho(t)|\mathbf{v}(t)\rangle=\langle\mathbf{v}(t)|e^{\gamma t}\rho_{0}|\mathbf{v}(t)\rangle=|c_{+}|^{2}+|c_{-}|^{2}=1$ (where we use $\langle u_{m}|\rho_{0}|u_{n}\rangle=\delta_{mn}$).
$\frac{d}{dt}\langle\mathbf{v}(t)|\rho(t)|\mathbf{v}(t)\rangle=\left(\frac{d}{dt}\langle\mathbf{v}(t)|\right)\rho|\mathbf{v}(t)\rangle+\langle\mathbf{v}(t)|\left(\frac{d}{dt}\rho\right)|\mathbf{v}(t)\rangle+\langle\mathbf{v}(t)|\rho\left(\frac{d}{dt}|\mathbf{v}(t)\rangle\right)$ (90)
$=\langle\mathbf{v}|\mathbf{L}_{\text{eff}}^{\dagger}\rho|\mathbf{v}\rangle+\langle\mathbf{v}|\dot{\rho}|\mathbf{v}\rangle+\langle\mathbf{v}|\rho\mathbf{L}_{\text{eff}}|\mathbf{v}\rangle$
$=\langle\mathbf{v}|\left[(iH_{PT}^{\dagger}-\frac{\gamma}{2})\rho+\gamma\rho+\rho(-iH_{PT}-\frac{\gamma}{2})\right]|\mathbf{v}\rangle=0.$
The non-orthogonality cause the complex transient dynamics with amplified noise and fluctuations at short time in both the cases of $s<0$ and s>0s>0[12, 13].
For $s<0$, in which case the eigenvalues relative to the global decay $i\gamma/2$ are imaginary and thus corresponds to exponential growth and decay, $\zeta_{3}(t)=C_{1}\cosh(\sqrt{-s}t)+C_{2}\sinh(\sqrt{-s}t)+C_{3}\sim e^{\sqrt{-s}t}$, $A(t)\sim-|\vec{\lambda}|^{2}\zeta_{3}(t)\sim-e^{2\Gamma t}$. If the system enters the PT-broken phase where $\zeta_{3}(t)$ grows exponentially, $A(t)$ will evolve accordingly to ensure that the norm of the fluctuation vector $\langle\mathbf{v}(t)|\rho(t)|\mathbf{v}(t)\rangle=1$, which satisfies the time-dependent Schrodinger equation $i\partial_{t}|\mathbf{v}(t)\rangle=i\mathbf{L}_{\text{eff}}|\mathbf{v}(t)\rangle$. For $s<0$ we further have,
$i\mathbf{L}_{\text{eff}}|u_{+}\rangle=\left(-\sqrt{(\Delta+2U|\alpha|^{2}-U^{2})^{2}-U^{2}|\alpha|^{4}}-i\frac{\gamma}{2}\right)|u_{+}\rangle=\left(i\frac{\sqrt{-s}}{2}-i\frac{\gamma}{2}\right)|u_{+}\rangle,$ (91)
$i\mathbf{L}_{\text{eff}}|u_{-}\rangle=\left(\sqrt{(\Delta+2U|\alpha|^{2}-Ur)^{2}-U^{2}|\alpha|^{4}}-i\frac{\gamma}{2}\right)|u_{-}\rangle=\left(-i\frac{\sqrt{-s}}{2}-i\frac{\gamma}{2}\right)|u_{-}\rangle,$
$-\langle u_{+}|i\mathbf{L}_{\text{eff}}^{\dagger}=\left(-i\frac{\sqrt{-s}}{2}+i\frac{\gamma}{2}\right)\langle u_{+}|,$
$-\langle u_{-}|i\mathbf{L}_{\text{eff}}^{\dagger}=\left(i\frac{\sqrt{-s}}{2}+i\frac{\gamma}{2}\right)\langle u_{-}|$
The dynamics of metric elements read
$\langle u_{+}|\dot{\rho}(t)|u_{+}\rangle$ $=-\langle u_{+}|(\mathbf{L}_{\text{eff}}^{\dagger}\rho+\rho\mathbf{L}_{\text{eff}})|u_{+}\rangle=-2(\frac{\sqrt{-s}}{2}-\gamma/2)\langle u_{+}|\rho(t)|u_{+}\rangle\implies\langle u_{+}|\rho(t)|u_{+}\rangle=\langle u_{+}|\rho(0)|u_{+}\rangle e^{-(2\frac{\sqrt{-s}}{2}-\gamma)t},$ (92)
$\langle u_{-}|\dot{\rho}(t)|u_{-}\rangle$ $=-\langle u_{-}|(\mathbf{L}_{\text{eff}}^{\dagger}\rho+\rho\mathbf{L}_{\text{eff}})|u_{-}\rangle=-2(-\frac{\sqrt{-s}}{2}-\gamma/2)\langle u_{-}|\rho(t)|u_{-}\rangle\implies\langle u_{-}|\rho(t)|u_{-}\rangle=\langle u_{-}|\rho(0)|u_{-}\rangle e^{(2\frac{\sqrt{-s}}{2}+\gamma)t},$
$\langle u_{+}|\dot{\rho}(t)|u_{-}\rangle$ $=-\langle u_{+}|(\mathbf{L}_{\text{eff}}^{\dagger}\rho+\rho\mathbf{L}_{\text{eff}})|u_{-}\rangle=-[(\frac{\sqrt{-s}}{2}-\gamma/2)+(-\frac{\sqrt{-s}}{2}-\gamma/2)]\langle u_{+}|\rho(t)|u_{-}\rangle=\gamma\langle u_{+}|\rho(t)|u_{-}\rangle\implies\langle u_{+}|\rho(t)|u_{-}\rangle=\langle u_{+}|\rho(0)|u_{-}\rangle e^{\gamma t}$
The norm reads
$\langle\mathbf{v}(t)|\rho(t)|\mathbf{v}(t)\rangle$ $=e^{-\gamma t}\left(c_{+}^{*}e^{\Gamma t}\langle u_{+}|+c_{-}^{*}e^{-\Gamma t}\langle u_{-}|\right)\rho(t)\left(c_{+}e^{\Gamma t}|u_{+}\rangle+c_{-}e^{-\Gamma t}|u_{-}\rangle\right)$ (93)
$=e^{-\gamma t}[|c_{+}|^{2}e^{2\Gamma t}R_{++}(t)+|c_{-}|^{2}e^{-2\Gamma t}R_{--}(t)+c_{+}^{*}c_{-}R_{+-}(t)+c_{-}^{*}c_{+}R_{-+}(t)]$
$=|c_{+}|^{2}R_{++}(0)+|c_{-}|^{2}R_{--}(0)+c_{+}^{*}c_{-}R_{+-}(0)+c_{-}^{*}c_{+}R_{-+}(0)=\langle\mathbf{v}(0)|\rho(0)|\mathbf{v}(0)\rangle=1,$
where the initial state is normalized under the initial metric such that $\langle\mathbf{v}(t)|\rho(t)|\mathbf{v}(t)\rangle=\langle\mathbf{v}(0)|\rho(0)|\mathbf{v}(0)\rangle=1$.
The Euclidean norm $\langle\mathbf{v}(t)|\mathbf{v}(t)\rangle=|\delta a|^{2}+|\delta a^{\dagger}|^{2}$ growing exponentially toward infinity, representing the instability due to the drive, making the probability non-conserved. The metric norm reads $\langle\mathbf{v}(t)|\rho_{0}|\mathbf{v}(t)\rangle=e^{-\gamma t}(|c_{+}|^{2}e^{2\Gamma t}+|c_{-}|^{2}e^{-2\Gamma t})$ which grows exponentially as $e^{(2\Gamma-\gamma)t}$. $\rho(t)$ is now no longer $e^{\gamma t}\rho_{0}$ with static $\rho_{0}$. It follows $\rho(t)=e^{\gamma t}\rho_{core}(t)$, where $\rho_{core}$ evolves to cancel $e^{\pm 2\Gamma t}$, such that $\langle\mathbf{v}(t)|\rho(t)|\mathbf{v}(t)\rangle=1$. The unitarity restoration condition $i\mathbf{L}_{\text{eff}}^{{\dagger}}\rho(t)-\rho(t)i\mathbf{L}_{\text{eff}}=i\dot{\rho}(t)$ ensures that the metric $\rho(t)$ evolves against the non-Hermitian Langevin drift of $\mathbf{L}_{\text{eff}}$ to keep the norm of the fluctuations invariant, i.e., $\frac{d}{dt}\langle\mathbf{v}(t)|\rho(t)|\mathbf{v}(t)\rangle=0$, and thus $\langle\mathbf{v}(t)|\mathbf{L}_{\text{eff}}^{\dagger}\rho(t)|\mathbf{v}(t)\rangle+\langle\mathbf{v}(t)|\dot{\rho}(t)|\mathbf{v}(t)\rangle+\langle\mathbf{v}(t)|\rho(t)\mathbf{L}_{\text{eff}}|\mathbf{v}(t)\rangle=0$ due to the drift equations $\dot{\mathbf{v}}(t)=\mathbf{L}_{\text{eff}}\mathbf{v}(t)$ and $\langle\dot{\mathbf{v}}(t)|=\langle\mathbf{v}(t)|\mathbf{L}_{\text{eff}}^{\dagger}$. Because a matrix with complex eigenvalues cannot meet the quasi-Hermiticity condition. In the PT broken phase, the system is inherently non-conservative (gaining/losing energy exponentially). The metric $\rho(t)$ serves as a dynamical renormalization tool and ensures the conserved probability density (unity metric norm).
The geometric core part satisfies the evolution based on the non-dissipative part $H_{PT}$
$i\dot{\rho}_{core}=H_{PT}^{\dagger}\rho_{core}-\rho_{core}H_{PT},$ (94)
and thus $H_{PT}^{\dagger}\rho(t)-\rho(t)H_{PT}=e^{\gamma t}(H_{PT}^{\dagger}\rho_{core}-\rho_{core}H_{PT})=e^{\gamma t}i\dot{\rho}_{core}$. In the $\mathcal{PT}$-symmetric phase, $\rho_{core}=\rho_{0}$ is static, so $i\dot{\rho}_{0}=0$. In the $\mathcal{PT}$-broken phase, $\rho_{core}(t)$ must evolve to compensate for complex eigenvalues. For the full scaling metric $\rho(t)=e^{\gamma t}\rho_{core}(t)$, the evolution under the full Hamiltonian $H$ is
$i\dot{\rho}(t)=(H_{PT}^{\dagger}\rho(t)-\rho(t)H_{PT})+i\gamma\rho(t)$ (95)
Substituting $\rho(t)=e^{\gamma t}\rho_{core}(t)$, we find
$i(\gamma e^{\gamma t}\rho_{core}(t)+e^{\gamma t}\dot{\rho}_{core}(t))=e^{\gamma t}(H_{PT}^{\dagger}\rho_{core}(t)-\rho_{core}(t)H_{PT})+i\gamma e^{\gamma t}\rho_{core}(t)$ (96)
This confirms that the global scaling $e^{\gamma t}$ handles the identity-like dissipation $i\gamma/2$, while $\rho_{core}(t)$ handles the internal $\mathcal{PT}$ geometry. Using the Dyson map $\rho(t)^{1/2}=e^{\gamma t/2}\rho_{core}^{1/2}(t)$, the mapped Hermitian operator reads
$e^{\gamma t/2}\rho_{core}^{1/2}(H_{PT}-i\frac{\gamma}{2})e^{-\gamma t/2}\rho_{core}^{-1/2}+i(\frac{\gamma}{2}e^{\gamma t/2}\rho_{core}^{1/2}+e^{\gamma t/2}\dot{\rho}_{core}^{1/2})(e^{-\gamma t/2}\rho_{core}^{-1/2})=\rho_{core}^{1/2}(t)H_{PT}\rho_{core}^{-1/2}(t)+i\dot{\rho}_{core}^{1/2}(t)\rho_{core}^{-1/2}(t)$ (97)
For $s<0$, the eigenvalues of $i\mathbf{L}_{\text{eff}}$ is $\lambda_{\pm}=-i\gamma/2\pm i\sqrt{U^{2}|\alpha|^{4}-(\Delta+2U|\alpha|^{2}-U)^{2}}$ and the corresponding eigenvectors are $|u_{\pm}\rangle$. Then the fluctuation vector $|\mathbf{v}(t)\rangle$ evolves as a superposition of these two eigenvectors $|\mathbf{v}(t)\rangle=e^{(-\gamma/2)t}(c_{+}e^{(\sqrt{-s}/2)t}|u_{+}\rangle+c_{-}e^{(-\sqrt{-s}/2)t}|u_{-}\rangle$, where $c_{\pm}$ are expansion coefficients determined by initial conditions at $t=0$ ($|\mathbf{v}(0)\rangle=c_{+}|u_{+}\rangle+c_{-}|u_{-}\rangle$). In non-Hermitian systems, these two eigenvectors are generally non-orthogonal. The Euclidean norm behaves as
$\langle\mathbf{v}(t)|\mathbf{v}(t)\rangle=e^{-\gamma t}\left[|c_{+}|^{2}e^{2(\sqrt{-s}/2)t}\langle u_{+}|u_{+}\rangle+|c_{-}|^{2}e^{-2(\sqrt{-s}/2)t}\langle u_{-}|u_{-}\rangle+2\text{Re}(c_{+}^{*}c_{-}\langle u_{+}|u_{-}\rangle)\right]$ (98)
which $\sim e^{(-\gamma+\sqrt{-s})t}$ at long time. The individual instantaneous eigenvectors $|u_{\pm}\rangle$ with complex eigenvalues do not maintain a simple biorthogonal relationship under a single evolving metric.
In Hermitian limit ($\gamma\to 0$ and $U\to 0$ in the relevant terms), the metric $\rho(t)$ and Dyson map tend to the identity operator. As the non-Hermitian contribution to the Hamiltonian vanishes, the right eigenvectors $|u_{\pm}\rangle$ become orthogonal ($\langle u_{+}|u_{-}\rangle\to 0$), and since the left and right eigenvectors coincide ($|l_{n}\rangle=|u_{n}\rangle$), the metric simplifies to $\rho=\sum_{n}|u_{n}\rangle\langle u_{n}|=\mathbb{I}$.
IX Appendix: Green’s Functions and Liouvillian Spectrum
In the limit of strong driving and dissipation, the standard Hamiltonian formalism is insufficient to describe the collective response. We treat the lattice problem by first solving the single-site effective Hamiltonians self-consistently coupled to their neighbors, governed by the non-Hermitian Liouvillian superoperator $\mathcal{L}$. The many-body renormalization effects are then incorporated via a diagrammatic T-matrix expansion base on the NESS.
In the mean-field approximation, the lattice problem is decoupled into single-site effective Hamiltonians self-consistently coupled to their neighbors. Unlike equilibrium systems governed by Hermitian Hamiltonians, the local physics here is governed by the non-Hermitian Liouvillian superoperator $\mathcal{L}$.
The linear response of the system to a weak probe field is encoded in the local retarded Green’s function. In the frequency domain, this can be expressed using the resolvent of the effective Liouvillian superoperator
$G^{R}_{loc}(\omega)$ $=\text{Tr}\left[\hat{a}\frac{1}{-i\omega-\mathcal{L}_{eff}}(\hat{a}^{\dagger}\hat{\rho}_{SS})\right]-\text{Tr}\left[\hat{a}^{\dagger}\frac{1}{-i\omega-\mathcal{L}_{eff}}(\hat{a}\hat{\rho}_{SS})\right]^{\dagger}.$ (99)
The effective Liouvillian superoperator $\mathcal{L}_{eff}$, which governs the time evolution of fluctuations $\delta\hat{\rho}$ linearized around the NESS density matrix $\hat{\rho}_{SS}$. By utilizing the spectral decomposition of $\mathcal{L}_{eff}=\sum_{\alpha}\lambda_{\alpha}|\breve{\rho}_{\alpha}\rangle\langle\breve{O}_{\alpha}|$, where $\lambda_{\alpha}=\Gamma_{\alpha}-i\omega_{\alpha}$ are the complex eigenvalues (with $\Gamma_{\alpha}\leq 0$ for stability), the Green’s function is
$G^{R}_{loc}(\omega)=\sum_{\alpha}\left[\frac{\text{Tr}(\hat{a}\breve{\rho}_{\alpha})\text{Tr}(\breve{O}_{\alpha}\hat{a}^{\dagger}\hat{\rho}_{SS})}{-i\omega-\lambda_{\alpha}}-\frac{\text{Tr}(\hat{a}^{\dagger}\breve{\rho}_{\alpha})^{*}\text{Tr}(\breve{O}_{\alpha}\hat{a}\hat{\rho}_{SS})^{*}}{i\omega-\lambda_{\alpha}^{*}}\right].$ (100)
The poles of $G^{R}_{loc}(\omega)$ identify the elementary excitations of the NESS. and correspond to the complex eigenvalues $\lambda_{\alpha}=\Gamma_{\alpha}-i\omega_{\alpha}$ of $\mathcal{L}_{eff}$.
The spectral function (local density of states) is defined as $\rho(\omega)=-\frac{1}{\pi}\text{Im}\left[G^{R}_{loc}(\omega)\right]$. For the Bose-Hubbard model with Kerr nonlinearity $U$ and drive $F$, the spectral function exhibits distinct features depending on the ratio $F/\gamma$. The spectral function $\rho(\omega)=-\frac{1}{\pi}\text{Im}[G^{R}_{loc}(\omega)]$ characterizes the energy distribution of different modes.
In linear regime with weak drive $F\ll\gamma,U$, the system remains close to the vacuum state. The non-linearity is negligible, and the Liouvillian spectrum is dominated by the single-particle loss $\gamma\mathcal{D}[\hat{a}$], leading to a single pole at $\lambda=-\gamma/2-i\Delta$. This yields the standard Lorentzian spectral function with peak centered at the detuning $\Delta$,
$\rho(\omega)\approx\frac{1}{\pi}\frac{\gamma/2}{(\omega-\Delta)^{2}+(\gamma/2)^{2}}.$ (101)
In non-linear regime with strong drive $F\gg\gamma$, as the drive increases, the photon number $n$ grows, and the interaction term $U\hat{n}(\hat{n}-1)$ becomes significant. The system behaves like a driven anharmonic oscillator. The excitation spectrum splits due to the dressing of energy levels by the drive field, leading to a multi-peak structure analogous to the mollow triplet.
As the coherent amplitude $\langle\hat{a}\rangle$ increases, the Kerr term $U\hat{n}(\hat{n}-1)$ induces energy-level transitions between dressed states. The Liouvillian spectrum splits into multiple branches. The resulting spectral function exhibits a mollow-triplet-like structure, where the spectral function can be approximated as a sum of resonances,
$\rho(\omega)\approx\sum_{j\in\{-1,0,1\}}\frac{A_{j}\Gamma_{j}}{(\omega-[\tilde{\omega}_{0}+j\Omega_{R}])^{2}+\Gamma_{j}^{2}},$ (102)
where $\tilde{\omega}_{0}=\Delta+2U\langle n\rangle$ is the Stark-shifted central frequency reflects the Stark-shifted resonance and $\Omega_{R}\approx 2F\sqrt{\langle n\rangle}$ is the vacuum Rabi splitting (effective Rabi frequency) induced by the drive acting on the non-linear medium.
The spatial coupling $J$ is incorporated by treating the hopping term as a perturbation on the local NESS. To analyze the stability of the spatial structure, we use the random phase approximation (RPA), and assume the fluctuations at different sites are coupled only via the mean-field dispersion. The lattice retarded Green’s function is related to the local Green’s function via the Dyson equation
$[G_{latt}^{R}(k,\omega)]^{-1}=[G^{(0)}_{loc}(\omega)]^{-1}-\epsilon_{k},$ (103)
where $\epsilon_{k}=-2J\cos(k)$ is the lattice dispersion relation (in 1D). This formulation maps the local Liouvillian poles onto collective bands in $k$-space. The dispersion of these collective modes is determined by the condition $\text{Re}[G_{loc}^{-1}(\omega)]=\epsilon_{k}$, while the imaginary part of $G_{loc}^{-1}$ provides the damping across the Brillouin zone. Note that $\epsilon_{k}$ is simply the momentum-space representation (Fourier transform) of the nearest-neighbor hopping term, necessary here to analytically diagonalize the spatial degrees of freedom.
Beyond the coherent propagation described by RPA, we account for the residual scattering between collective excitations. Adapting the T-matrix (ladder) approximation from polaron theory, we define the scattering of two excitations with total momentum $Q$ and total frequency $\Omega$. The non-equilibrium pair propagator (driven polarization bubble) $\Pi(Q,\Omega)$ can be obtained by the lattice Green’s functions. In the frequency domain, this involves a convolution that accounts for the dissipative character of NESS
$\Pi(Q,\Omega)=\frac{i}{2\pi}\int d\omega^{\prime}\frac{1}{L}\sum_{k}G_{latt}^{R}(k,\omega^{\prime})G_{latt}^{R}(Q-k,\Omega-\omega^{\prime}).$ (104)
Unlike equilibrium systems, the integration path is determined by the analytic structure of $G_{latt}^{R}$ in complex plane, where poles reside in the lower half-plane due to $\Gamma_{\alpha}<0$. The interaction is dressed by multiple scattering events to form the renormalized T-matrix $\mathcal{T}(Q,\Omega)$:
$\mathcal{T}(Q,\Omega)=\frac{U}{1-U\Pi(Q,\Omega)}.$ (105)
Divergences in $\mathcal{T}(Q,\Omega)$ signal the formation of repulsively bound pairs (doublons) in the lattice. In the driven-dissipative case, the condition $1=U\text{Re}[\Pi(Q,\Omega)]$ determines the energy of these pairs, while $\text{Im}[\Pi]$ determines their finite lifetime. The final renormalization of the single-particle excitation spectrum is given by the ladder self-energy $\Sigma_{ladder}$:
$\Sigma_{ladder}(k,\omega)=\int\frac{d\Omega}{2\pi}\frac{1}{L}\sum_{q}G_{latt}^{R}(q-k,\Omega-\omega)\mathcal{T}(q,\Omega).$ (106)
The total lattice response is then governed by the fully renormalized Green’s function $\tilde{G}(k,\omega)=[G_{latt}^{R}(k,\omega)^{-1}-\Sigma_{ladder}(k,\omega)]^{-1}$. A key consequence of this calculation is the modification of the spectral weight $Z_{k}$ and the introduction of collision-induced broadening. If $\text{Im}[\Sigma_{ladder}]$ exceeds the drive-induced gain in specific momentum sectors, the spatial NESS becomes unstable, leading to a breakdown of the homogeneous phase.